### v1

# Modular analogs of character formulas and multiplicities of induced representations

Abstract.
Let $G$ be a finite group, $k$ a characteristic-$p$ field, and $\pi$ be a simple $k[G]$-module. We show that if $\pi$ is in a $p$-nilpotent block with cyclic defect group, then $\pi$ appears with multiplicity at most one in the socle and cosocle of induced $k[G]$-modules that are multiplicity-free in characteristic zero and satisfy some additional hypotheses. We show that this applies, in particular, to certain Deligne–Lusztig modules $\pi$. As an application, we prove that well-known character formulas for cuspidal representations of $\mathrm{GL}_{2}(\mathbb{F}_{\ell})$ have modular analogs as statements about short exact sequences of lattices. We also prove modular analogs of character formulas for representations coming from higher-depth supercuspidal representations of $\mathrm{GL}_{2}$. The appendix by the authors and R. Zhang is an erratum to “Modular Gelfand pairs and multiplicity-free representations” by R. Zhang, which states some multiplicity-freeness results in overly broad generality; we provide counterexamples and partial corrections to these claims.
1. Introduction
Many interesting irreducible characteristic-zero representations of finite groups are best described by character formulas. Typically, these formulas involve writing the character of the representation as a linear combination of characters of representations that are better understood (for instance, those that are induced from subgroups that are simpler). Since the category of characteristic-zero representations is semisimple, the character formula is essentially enough to completely understand the representation. However, if one wants to understand the modular theory (that is, over characteristic-$p$ fields), then the character formula can be insufficient. Indeed, the impetus for this paper was an attempt to compute the cohomology of the mod-$p$ reduction of a cuspidal representation of $\operatorname{GL}_{2}(\mathbb{F}_{\ell})$, where the character formula is well known, but not particularly useful when computing cohomology. In this paper, we consider the question of finding modular analogs of character formulas, and describe a situation where this question comes down to a question about multiplicity-freeness, and prove some results relating multiplicity-freeness in characteristic zero and characteristic $p$.
The following is an example that illustrates our main results in the case of the group $G=\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ for a prime number $\ell$. We will use some standard notation here without explaining it; see Section 5 for a complete explanation. Let $T\leq G$ be a non-split torus, $\theta$ a primitive character of $T$, and $\pi_{\theta}$ be the cuspidal representation of $G$ associated to $\theta$. There is a well-known formula for the character of $\pi_{\theta}$ that comes from an equality of virtual representations
$\pi_{\theta}=\mathrm{Ind}_{ZU}^{G}\psi-\mathrm{Ind}_{T}^{G}\theta.$
For a prime number $p$ with $\ell\equiv-1\pmod{p}$, we prove that this equality has an analog for $\mathbb{Z}_{p}$-lattices in these representations. That is, there is an exact sequence of lattices
(1.1) $0\to\tilde{\pi}_{\theta}\to\tilde{e}\mathrm{Ind}_{ZU}^{G}\psi\to\tilde{e}\mathrm{Ind}_{T}^{G}\theta\to 0$
where $\tilde{\pi}_{\theta}$ is a lattice in $\pi_{\theta}$, and where $\tilde{e}$ is the block idempotent corresponding to $\pi_{\theta}$. This is useful, for instance, because it allows one to compute the Tate cohomology $\hat{H}^{i}(G^{\prime},\tilde{\pi}_{\theta})$ for all subgroups $G^{\prime}\leq G$ (see Theorem 1.5). In Section 1.4, we discuss an application of this kind of computation to questions of finding minimal and non-minimal lifts of mod-$p$ modular forms.
1.1. Modular analogs
We now explain what we mean by character formulas and their modular analogs. We fix some notation now and for the entire paper: $p$ is a prime number, $G$ is a finite group, and $K$ is a $p$-adic local field that contains the $|G|$th-roots of unity, with valuation ring $\mathcal{O}$, uniformizer $\varpi$, and residue field $k=\mathcal{O}/\varpi\mathcal{O}$.
We will also use the following notation for the introduction. For a finitely-generated $K[G]$-module $M$, we will let $M_{\mathcal{O}}$ denote a choice of $\mathcal{O}$-lattice in $M$ and let ${\overline{M}}_{\mathcal{O}}$ denote $M_{\mathcal{O}}\otimes_{\mathcal{O}}k$. Let $\pi$ be an irreducible $K[G]$-module and suppose there is an equality
(1.2) $\pi=M-M^{\prime}$
of virtual representations, where $M$ and $M^{\prime}$ are $K[G]$-modules. We refer to (1.2) as a character formula for $\pi$. Then a modular analog of (1.2) is a short exact sequence of $\mathcal{O}[G]$-modules
(1.3) $0\to\pi_{\mathcal{O}}\to M_{\mathcal{O}}\to M^{\prime}_{\mathcal{O}}\to 0.$
Note that the existence of the sequence (1.3) implies the character formula (1.2), but not the other way around in general. However, we do have the following simple criterion.
Lemma 1.1.
Assume that $\pi=M-M^{\prime}$ as virtual representations. Suppose that $M$ has a lattice $M_{\mathcal{O}}$ that is a projective $\mathcal{O}[G]$-module and that $M^{\prime}$ has a lattice $M^{\prime}_{\mathcal{O}}$ such that, for every simple $k[G]$-module $\eta$,
(1.4) $\dim_{k}\mathrm{Hom}_{k[G]}({\overline{M}}_{\mathcal{O}},\eta)\geq\dim_{k}\mathrm{Hom}_{k[G]}({\overline{M^{\prime}}}_{\mathcal{O}},\eta).$
Then there is a lattice $\pi_{\mathcal{O}}$ in $\pi$ such that there is a short exact sequence of $\mathcal{O}[G]$-modules
$0\to\pi_{\mathcal{O}}\to M_{\mathcal{O}}\to M^{\prime}_{\mathcal{O}}\to 0.$
Proof.
The assumption (1.4) implies that there is a surjective $k[G]$-module homomorphism ${\overline{f}}:\mathrm{cosoc}({\overline{M}}_{\mathcal{O}})\twoheadrightarrow\mathrm{cosoc}({\overline{M^{\prime}}}_{\mathcal{O}})$. Since $M_{\mathcal{O}}$ is projective, there is $\mathcal{O}[G]$-module homomorphism $f:M_{\mathcal{O}}\to M^{\prime}_{\mathcal{O}}$ such that $\mathrm{cosoc}(f\otimes_{\mathcal{O}}k)={\overline{f}}$. By Nakayama’s lemma, $f$ is surjective. Since $\pi=M-M^{\prime}$, the kernel of $f$ is a lattice in $\pi$. ∎
Abstract character formulas and their modular analogs are not very interesting: for instance, one could take $M=\pi$ and $M^{\prime}=0$ in (1.2), and the character formula is true, but contains no new information. To be of use, character formulas must involve modules $M$ and $M^{\prime}$ that are simpler or better understood than $\pi$. The same is true for modular analogs and the lattices $M_{\mathcal{O}}$ and $M^{\prime}_{\mathcal{O}}$. In this paper, we are interested in character formulas that involve modules $M$ and $M^{\prime}$ that are induced from proper subgroups $H,H^{\prime}\leq G$, and we will require that the lattices $M_{\mathcal{O}}$ and $M^{\prime}_{\mathcal{O}}$ be induced from the same subgroups $H$ and $H^{\prime}$ [^1] To be more precise, we actually allow lattices that are blocks in such inductions.. To put ourselves in the setting of the lemma, we will assume that $p\nmid\#H$. Note that the two sides of (1.4) are the multiplicities of $\eta$ in the cosocles of of ${\overline{M}}_{\mathcal{O}}$ and ${\overline{M^{\prime}}}_{\mathcal{O}}$, so we are led to the problem of computing multiplicities of representations in cosocles of induced representations. For the remainder of the introduction, we will abuse notation by writing $\overline{M}$} instead of ${\overline{M}}_{\mathcal{O}}$.
1.2. Multiplicities of modular inductions and commutativity of their endomorphism rings
Schur’s lemma implies that a semisimple module is multiplicity free if and only if its endomorphism ring is commutative. This is a useful tool for showing that a module is multiplicity free. In characteristic $p$, this tool cannot always be used because neither implication is true without the assumption that $\overline{M}$} is semisimple. Nonetheless, knowing the commutativity of $\mathrm{End}({\overline{M}})$ can still be useful. Indeed, if one also knows that the map $\mathrm{End}({\overline{M}})\to\mathrm{End}(\mathrm{cosoc}({\overline{M}}))$ is surjective, then one can deduce that $\mathrm{End}(\mathrm{cosoc}({\overline{M}}))$ is commutative and hence, since $\mathrm{cosoc}({\overline{M}})$ is semisimple, that $\mathrm{cosoc}({\overline{M}})$ is multiplicity free. This is the approach taken in [Zha24], where one of the main results is the claim that, if an induced representation of a finite group is multiplicity free in characteristic zero, then the cosocle of its mod-$p$ reduction is also multiplicity free. Unfortunately, [Zha24] contains errors, and this claim is incorrect in the generality stated.
The appendix to this paper is an erratum to [Zha24], and contains counterexamples and partial corrections to its main results. The first main step in [Zha24, Corollary 4.8] is to show that, if $M$ is an induced representation and $\mathrm{End}(M)$ is commutative, then $\mathrm{End}({\overline{M}})$ is commutative. This is not true in general, but we give a correction in Proposition A.3 that is sufficient for our purposes. The second main step is to show that, if $\overline{M}$} is an induction of an irreducible representation, then the map $\mathrm{End}({\overline{M}})\to\mathrm{End}(\mathrm{cosoc}({\overline{M}}))$ is surjective. This is also false in general. We prove some criteria for when this holds, but they are not sufficient for our main cases of interest, so we require a new technique that allow us to prove, in the cases of interest, that $\mathrm{cosoc}({\overline{M}})$ is multiplicity free using the commutativity of $\mathrm{End}({\overline{M}})$.
1.3. Nilpotent blocks and Deligne–Lusztig theory
To prove the multiplicity-freeness of $\mathrm{cosoc}({\overline{M}})$, we use techniques of modular representation theory together with some simple algebraic observations. The first observation is that, if $R=k[x]/(x^{n})$ is a monogenic local $k$-algebra, then the well-known structure theory implies that an $R$-module with commutative endomorphism ring is cyclic, and that cyclic $R$-modules have multiplicity-free socle and cosocle. The second observation is that the same is true for rings $R$ that are Morita equivalent to a monogenic local $k$-algebra.
We can then appeal to a result of Broué and Puig from modular representation theory to see why these observations are relevant. Their result says that if $e$ is a $p$-nilpotent block in $k[G]$ that has cyclic defect group $D$, then $ek[G]$ is Morita-equivalent to $k[D]$ (which is a monogenic local ring) [BP80, pg. 120]. We show in Lemma 3.4 that Broué and Puig’s result implies the following.
Theorem 1.2.
Let $e$ be a $p$-nilpotent block in $k[G]$ that has cyclic defect group, and let $\pi$ be a simple $ek[G]$-module. If $\overline{M}$} is a finitely generated $k[G]$-module such that $\mathrm{End}_{k[G]}({\overline{M}})$ is commutative and $e{\overline{M}}\neq 0$, then
$\dim_{k}\mathrm{Hom}_{k[G]}({\overline{M}},\pi)=\dim_{k}\mathrm{Hom}_{k[G]}(\pi,{\overline{M}})=1.$
Using Deligne–Lusztig theory, we show that some irreducible representations of Chevalley groups in characteristic $\ell\neq p$ are in $p$-nilpotent blocks with cyclic defect group. Hence, the above theorem applies to these representations.
Theorem 1.3.
Let $F$ be a finite field of characteristic $\ell$. Let $G=\mathbf{G}(F)$ for a reductive group $\mathbf{G}$. Suppose that there is a maximal torus $\mathbf{T}\subset\mathbf{G}$ such that the $p$-Sylow subgroup of $T=\mathbf{T}(F)$ is cyclic and is a $p$-Sylow subgroup of $G$.
Let $\theta:T\to\mathbb{C}^{\times}$ be a character whose mod-$p$ reduction $\overline{\theta}$} is in general position. Then the Deligne–Lusztig irreducible representation $\pi_{\theta}$ associated to $\theta$ is in a $p$-nilpotent block with cyclic defect group.
This is proven in Theorem 4.3 below. Combining these results with the results in the appendix about commutativity of endomorphism rings of induced representations yields the following.
Theorem 1.4.
Let $G$, $\theta$ and $\pi_{\theta}$ be as in Theorem 1.3, and let ${\overline{\pi}}_{\theta}$ be the reduction of a lattice in $\pi_{\theta}$.
Let $H$ be a subgroup of $G$ and $\psi:H\to\mathbb{C}^{\times}$ be a character. For $s\in G$, let $H_{s}=H\cap s^{-1}Hs$ and let $\psi^{s}:H_{s}\to\mathbb{C}^{\times}$ be $\psi^{s}(x)=\psi(sxs^{-1})$. Assume that $\mathrm{Ind}_{H}^{G}\psi$ is multiplicity free, and that there is no $s\in G$ such that the character $\psi^{-1}\psi^{s}:H_{s}\to\mathbb{C}^{\times}$ is nontrivial and has $p$-power order. Then,
$\dim_{k}\mathrm{Hom}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\psi}},{\overline{\pi}}_{\theta})=\dim_{k}\mathrm{Hom}_{k[G]}({\overline{\pi}}_{\theta},\mathrm{Ind}_{H}^{G}{\overline{\psi}})=1,$
Proof.
The assumptions on $\psi$ imply that $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\psi}})$ is commutative by Proposition A.3, so the result follows by combining Theorem 1.2 and Theorem 1.3. ∎
Finally, we can combine this result with Lemma 1.1 to get a statement about modular analogs of character formulas.
Theorem 1.5.
Let $G$, $\theta$, $\pi_{\theta}$, $H$, and $\psi$ be as in Theorem 1.4. Suppose that there is a projective $\mathcal{O}[G]$-module $M_{\mathcal{O}}$ such that there is an equality of virtual representations
$\pi_{\theta}=M-\mathrm{Ind}_{H}^{G}\psi.$
Let $\tilde{e}\in\mathcal{O}[G]$ be the central idempotent for the block containing $\pi_{\theta}$. Then there is a lattice $\tilde{\pi}_{\theta}$ in $\pi_{\theta}$ and an exact sequence of $\mathcal{O}[G]$-modules
$0\to\tilde{\pi}_{\theta}\to\tilde{e}M_{\mathcal{O}}\to\tilde{e}\mathrm{Ind}_{H}^{G}\psi\to 0.$
In particular, for every subgroup $G^{\prime}\leq G$ and every $n\in\mathbb{Z}$ there is an isomorphism of Tate cohomology groups
$\hat{H}^{n}(G^{\prime},\tilde{\pi}_{\theta})\cong\hat{H}^{n-1}(G^{\prime},\tilde{e}\mathrm{Ind}_{H}^{G}\psi).$
Proof.
By Theorem 1.4, the inequality (1.4) holds, and the first statement follows by Lemma 1.1. The statement about Tate cohomology holds because the projective module $\tilde{e}M_{\mathcal{O}}$ is cohomologically trivial. ∎
This theorem implies the modular character formula (1.1) for cuspidal representations of $\operatorname{GL}_{2}(\mathbb{F}_{\ell})$, as we explain in Section 5 below. In the same section, we also prove similar results for finite-group representations that are relevant for the construction of higher-depth supercuspidal representations of $\ell$-adic $\operatorname{GL}_{2}$.
1.4. Application to modular forms and minimal lifts
We end this introduction with an application to modular forms and Galois representations that was the initial impetus for this project. Let $\mathrm{Gal}_{\mathbb{Q}}=\mathrm{Gal}({\overline{\mathbb{Q}}}/\mathbb{Q})$ and let ${\overline{\rho}}:\mathrm{Gal}_{\mathbb{Q}}\to\operatorname{GL}_{2}(k)$ be an absolutely irreducible two-dimensional mod-$p$ Galois representation. Serre [Ser87] formulated a precise conjecture about which $\overline{\rho}$} come from modular forms, and, when they do, what the minimal weight, level, and character of the form are. Later, he found that there was an error in the character part of his conjecture for $p\in\{2,3\}$ and gave a counterexample [Ser88, Contre-exemple, pg. 54] for $p=3$ using the modular form [LMF25, Newform orbit 13.2.e.a]. As Serre explains, the issue has to do with $3$-torsion in the cohomology of certain arithmetic subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$.
Theorem 1.5 can be used to study related questions. The minimal weight, level, and character are closely related to the notion of a minimally ramified lift of $\overline{\rho}$}. In fact, there is only one type of prime $\ell$, called vexing primes, for which the minimally ramified lift is not determined by the weight, level, and character; these are the primes $\ell\equiv-1\pmod{p}$ such that ${\overline{\rho}}|_{D_{\ell}}$ is irreducible and ${\overline{\rho}}|_{I_{\ell}}$ is irreducible, where $I_{\ell}\leq D_{\ell}\leq\mathrm{Gal}_{\mathbb{Q}}$ are an inertia group and decomposition group at $\ell$. For every vexing prime $\ell$ and eigenform $f$ associated to $\overline{\rho}$}, the local component at $\ell$ of the automorphic representation of $f$ is a supercuspidal representation whose Bernstein component is associated to some smooth character $\theta$ of $\mathbb{Z}_{\ell^{2}}^{\times}$. The minimally ramified (at $\ell$) lifts are the ones where the character $\theta$ has minimal order.
Example 1.6.
Let $p=3$ and $\ell=11$. Consider the newform $f$ of weight $2$ and level $\Gamma_{0}(121)$ with $q$-expansion
$f=q-q^{3}-2q^{4}-3q^{5}-2q^{9}+2q^{12}+\dots$
[LMF25, Newform orbit 121.2.a.b]. Using the Loeffler–Weinstein algorithm [LW12] in Sage [Sag25] reveals that the local component of $f$ at $11$ is a depth-zero supercuspidal representation associated to a character $\theta$ of $\mathbb{F}_{121}^{\times}$ that has order $4$. In particular, the reduction $\overline{\theta}$} modulo $3$ also has order $4$ and hence cannot factor through $\mathbb{F}_{11}^{\times}$, so it is in general position.
Let $\rho_{f}$ and $\overline{\rho}$} be the $3$-adic and mod-$3$ representations of $f$, respectively. The above calculation shows that $11$ is a vexing prime for $\overline{\rho}$} and that $\rho_{f}$ is a minimally ramified lift of $\overline{\rho}$}. Examining the list of newforms of weight $2$ and level $\Gamma_{1}(121)$ in [LMF25] quickly reveals that there are no other forms there congruent to $f$. In particular, in this case, there are no lifts of $\overline{\rho}$} of weight $2$ and level $\Gamma_{1}(121)$ that are not minimally ramified at $11$.
Theorem 1.5 can be used to explain this example and related phenomena. Let $\tilde{\pi}_{\theta}$ be an $\mathcal{O}$-lattice in the cuspidal representation of $\operatorname{GL}_{2}(\mathbb{F}_{11})$ associated to $\theta$. There is an exact sequence
$0\to H^{1}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta})\otimes_{\mathcal{O}}k\to H^{1}(\mathrm{SL}_{2}(\mathbb{Z}),{\overline{\pi}}_{\theta})\to H^{2}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta}),$
and $H^{2}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta})$ is an obstruction to finding a minimal lift. The theory of Farrell cohomology [Bro94, Corollary 7.4, pg. 293] implies that there is an isomorphism
$H^{2}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta})\cong H^{2}(P,\tilde{\pi}_{\theta})$
where $P\leq\mathrm{SL}_{2}(\mathbb{Z})$ is a subgroup of order $3$ (which is unique up to conjugacy). Hence, by Theorem 1.5,
$H^{2}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta})\cong H^{2}(P,\tilde{\pi}_{\theta})\cong H^{1}(P,e\mathrm{Ind}_{T}^{G}\theta).$
Using Mackey theory, it is easy to see that, as a $P$-module, $e\mathrm{Ind}_{T}^{G}\theta$ is isomorphic to $\theta|_{P}\oplus\theta^{c}|_{P}\oplus M$ for a projective $\mathcal{O}[P]$-module $M$. Hence,
$H^{2}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta})\cong H^{1}(P,\theta|_{P}\oplus\theta^{c}|_{P}).$
Since $\theta$ is minimal, it has order prime to $3$, so $\theta|_{P}=\theta^{c}|_{P}=\mathcal{O}$, and $H^{1}(P,\theta|_{P}\oplus\theta^{c}|_{P})=0$. Altogether, this says that there is no obstruction to finding a minimal lift.
However, if we had replaced $\theta$ with a non-minimal lift of $\overline{\theta}$}, then $\theta|_{P}$ and $\theta^{c}|_{P}$ are nontrivial, and the same computation would yield $H^{2}(\mathrm{SL}_{2}(\mathbb{Z}),\tilde{\pi}_{\theta})=H^{1}(P,\theta|_{P}\oplus\theta^{c}|_{P})\cong\mathbb{F}_{3}\oplus\mathbb{F}_{3}$. In other words, there is potentially an obstruction to finding non-minimal lifts, as observed in Example 1.6.
This circle of ideas will be explored more fully in future work.
1.5. Acknowledgements
P. W. thanks Bob Kottwitz, Jackie Lang and Rob Pollack for helpful conversations about modular representation theory and Deligne–Lusztig theory. R. Z. is extremely grateful to his coauthors for identifying and communicating the errors and for agreeing to work together on this note. He also thanks Karol Koziol for identifying some of the errors and pointing out Example A.5; Ofer Gabber for explaining the equivalence of self-projectivity and almost-projectivity when there are enough projectives; and Dipendra Prasad for pointing out Corollary A.16.
P. A. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2020-05915]. P. A. a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [numéro de référence RGPIN-2020-05915]. P. W. is supported by National Science Foundation CAREER Grant No. DMS-2337830. R. Z. is supported by National Science Foundation Grant No. DMS-2303280.
2. Multiplicity-freeness over monogenic local $k$-algebras
In this section, we consider finite-dimensional, unital, associative $k$-algebras $R$. For a (left) $R$-module $M$, we write $M^{*}=\mathrm{Hom}_{k}(M,k)$ for the dual $R$-module. We recall a situation where $\mathrm{End}_{R}(M)$ being commutative obvioulsy implies that the socle and cosocle of $M$ are multiplicity free: when $R$ is a monogenic local ring (or Morita-equivalent to a such a ring). We also recall the theory of blocks of $R$-modules.
2.1. Modules over monogenic local $k$-algebras
A $k$-algebra $S$ is called monogenic if $S$ is a quotient of $k[x]$, and called local if $S/\mathrm{Rad}(S)\cong k$; a finite-dimensional monogenic local $k$-algebra $S$ has the form $S=k[x]/(x^{m})$. For example, if $k$ has characteristic $p$ and $G$ is a cyclic $p$-group, then $k[G]$ is a monogenic local $k$-algebra. The following is an easy consequence of the structure theory for modules over monogenic local $k$-algebras.
Lemma 2.1.
Let $k$ be a field, let $S=k[x]/(x^{m})$ for some $m>0$, and let $R$ be a $k$-algebra such that there is a Morita equivalence of categories
$F:S\text{-mod}\to R\text{-mod}.$
Let $P=F(S)$ and let $\pi=F(k)$. Let $M$ be a finitely-generated nonzero $R$-module. The following are equivalent:
  1. (1)
    $\mathrm{cosoc}(M)\cong\pi$.
  2. (2)
    $M$ is indecomposable.
  3. (3)
    There is a surjection $P\twoheadrightarrow M$.
  4. (4)
    $\mathrm{End}_{R}(M)$ is commutative.
  5. (5)
    There is an injection $M\hookrightarrow P$.
  6. (6)
    $\mathrm{soc}(M)\cong\pi$.
Moveover, any of these are true for $M$ if and only if they are true for $M^{*}$.
Proof.
From the equivalence of categories, it suffices to prove this in the case $R=S$ with $P=S$ and $\pi=k$. First, the last statement is clear because (2) holds for $M$ if and only if it holds for $M^{*}$. Clearly (1) implies (2) and (3) implies (1), and, since every finitely-generated $S$-module is a direct sum of cyclic modules, (2) implies (3). If $M=S/I$ for an ideal $I$, then $\mathrm{End}_{S}(M)\cong S/I$ is commutative, so (3) implies (4). If $M=A\oplus B$ with $A,B\neq 0$, then the fact that $S$ is local implies that $\mathrm{Hom}_{S}(A,B)\neq 0$, and it follows easily that $\mathrm{End}_{S}(M)$ is not commutative, so (4) implies (2). This shows that (1)-(4) are equivalent, and are also equivalent to (1)-(4) for $M^{*}$. Since $S^{*}\cong S$ as $S$-modules, (5) for $M$ is equivalent to (3) for $M^{*}$, and since $\mathrm{soc}(M)^{*}=\mathrm{cosoc}(M^{*})$, (6) for $M$ is equivalent to (1) for $M^{*}$. ∎
The $k$-algebras $R$ that are Morita equivalent to $S=k[x]/(x^{m})$ are the $n\times n$-matrix algebras $R=\mathrm{M}_{n}(S)$, in which case the equivalence $F$ is given by $F(M)=M\otimes_{S}S^{n}$.
2.2. Blocks
The center $Z(R)$ of $R$ maps to the endomorphism ring of every $R$-module. A block in $R$ is a primitive idempotent $e$ in $Z(R)$. There is an isomorphism of rings $R\cong\prod_{e}eR$, where $e$ ranges over the blocks of $R$, that induces induces an isomorphism $M\cong\prod_{e}eM$ for every $R$-module $M$. For a simple $R$-module $\pi$, a block $e$ maps to an idempotent of $\mathrm{End}_{R}(\pi)$, which is either zero or one. We say that $\pi$ belongs to the block $e$ if $e$ acts as one on $\pi$. Every simple module belongs to exactly one block. If $\pi$ belongs to the block $e$, then every homomorphism $M\to\pi$ or $\pi\to M$ factors uniquely through $eM$.
Lemma 2.2.
Let $M$ be a finitely generated $R$-module such that $\mathrm{End}_{R}(M)$ is commutative, and let $e$ be a block of $R$ such that $eR$ is Morita equivalent to a monogenic local $k$-algebra. If $eM\neq 0$ and $\pi$ is a simple $R$-module that belongs to $e$, then
$\dim_{k}\mathrm{Hom}_{R}(\pi,M)=\dim_{k}\mathrm{Hom}_{R}(M,\pi)=1.$
Proof.
There is an isomorphism of rings, $End_{R}(M)\cong\prod_{e}\mathrm{End}_{eR}(eM)$, so $\mathrm{End}_{eR}(eM)$ is commutative. By Lemma 2.1,
$\mathrm{soc}(eM)\cong\mathrm{cosoc}(eM)\cong\pi.$
Since
$\mathrm{Hom}_{R}(\pi,M)=\mathrm{Hom}_{eR}(\pi,eM)=\mathrm{Hom}_{eR}(\pi,\mathrm{soc}(eM))$
and similarly $\mathrm{Hom}_{R}(M,\pi)=\mathrm{Hom}_{eR}(\mathrm{cosoc}(eM),\pi)$, this completes the proof. ∎
3. Applications of modular representation theory
In this section, we review the theory of blocks in modular representation theory. We recall a result of Broué and Puig [BP80] that says that nilpotent blocks with cyclic defect group are Morita equivalent to local monogenic algebras, which will allow us to apply the results of the previous section. Throughout this section, $G$ denotes a finite group.
3.1. Blocks in group algebras and defect groups
We recall Brauer’s theory of blocks, following [Isa06, Chapter 15] (see also [Bum15]). The reference [Isa06] works with representations over $\mathbb{C}$, everything holds without change for representations over a large enough field of characteristic zero (see [Ser77a, Chapter 12], for instance) like $K$, and we have translated all the results to the corresponding statements over $K$.
Let $\mathrm{Conj}(G)$ denote the set of conjugacy classes in $G$. The center $Z(K[G])$ of $K[G]$ is spanned by the indicator functions $1_{\mathcal{C}}$ for $\mathcal{C}\in\mathrm{Conj}(G)$. For an element or subset $S$ of $G$, let $C_{G}(S)$ be the centralizer of $S$ in $G$. Let $\mathrm{Irr}(G)$ denote the set of isomorphism classes of irreducible representations $\rho:G\to\operatorname{GL}_{n}(K)$ of $G$, and for $\rho\in\mathrm{Irr}(G)$, let $\chi_{\rho}:G\to K$ be its (trace) character.
For representations over $k$, a variant of character, called Brauer character, plays an important role. For a $k[G]$-module $\eta$ and an element $y\in G$ of prime-to-$p$ order, the eigenvalues of $\eta(y)$ in $k$, being roots of unity of prime-to-$p$ order, lift uniquely to elements of $\mathcal{O}^{\times}$ of the same order, and we let $\varphi_{\eta}(y)\in\mathcal{O}$ be the sum of these eigenvalues with multiplicity; the function $\varphi_{\eta}$ is called the Brauer character of $\eta$. The Brauer characters of simple $k[G]$-modules are called irreducible, and the set of them is denoted $\mathrm{IBr}(G)$; every Brauer character is a non-negative-integer combination of irreducible ones. If $\rho\in\mathrm{Irr}(G)$, then $\hat{\chi}_{\rho}$, the restriction of $\chi_{\rho}$ to the prime-to-$p$-order elements of $G$, is a Brauer character ([Isa06, Theorem 15.6, pg. 265]), so it can be written as
$\hat{\chi}_{\rho}=\sum_{\varphi\in\mathrm{IBr}(G)}d_{\rho,\varphi}\varphi,$
for nonnegative integers $d_{\rho,\varphi}$ called the decomposition numbers of $\rho$.
Irreducible representations of $G$ over $K$ come in blocks according to their reductions to $k$. Define the Brauer graph of $G$ to have vertex set $\mathrm{Irr}(G)$ and an edge connecting $\rho,\rho^{\prime}\in\mathrm{Irr}(G)$ if and only if there is $\varphi\in\mathrm{IBr}(G)$ such that $d_{\rho,\varphi}$ and $d_{\rho^{\prime},\varphi}$ are both nonzero. The set of connected components of the Brauer graph is denoted $\mathrm{Bl}(G)$ and its elements (thought of as subsets of $\mathrm{Irr}(G)$) are called blocks of representations. (This differs from the definition given in [Isa06, Definition 15.17 pg. 271] but is equivalent by [Isa06, Theorem 15.27, pg. 275].) For $B\in\mathrm{Bl}(G)$ and $\varphi\in\mathrm{IBr}(G)$, we say that $\varphi$ belongs to the block $B$ if there is $\rho\in B$ such that $d_{\rho,\varphi}\neq 0$. Each $\varphi$ belongs to a unique block ([Isa06, Theorem 15.19,pg. 272]).
Each $B\in\mathrm{Bl}(G)$ defines central idempotents $\tilde{e}_{B}\in\mathcal{O}[G]$ and $e_{B}\in k[G]$, as we now explain; this defines a bijection between $\mathrm{Bl}(G)$ and the set of blocks of $k[G]$ as in Section 2.2, justifying the similar notation and allowing us to use the two meanings of ‘block’ interchangeably. For $\rho\in\mathrm{Irr}(G)$, let $e_{\rho}\in Z(K[G])$ be the element
$e_{\rho}=\frac{\dim_{\mathbb{C}}\rho}{\#G}\sum_{g\in G}\chi_{\rho}(g)[g^{-1}].$
These $e_{\rho}$ are mutually orthogonal idempotents in $Z(K[G])$, so a sum of them over $\rho$ in a subset of $\mathrm{Irr}(G)$ is also an idempotent. Due to the #G\#G in the denominator, $e_{\rho}$ will not typically be in $\mathcal{O}[G]$. However, if $B\in\mathrm{Bl}(G)$, then the sum $\tilde{e}_{B}=\sum_{\rho\in B}e_{\rho}$ is in $\mathcal{O}[G]$, and, in fact, a nonempty subset $B^{\prime}\subset\mathrm{Irr}(G)$ is a union of blocks of representations if and only if $\sum_{\rho\in B^{\prime}}e_{\rho}\in\mathcal{O}[G]$ ([Isa06, Theorems 15.26, 15.27, pg. 275]). If $B\in\mathrm{Bl}(G)$, then $\tilde{e}_{B}$ defines a central idempotent $e_{B}$ in $k[G]$ by reduction; these $e_{B}$ are blocks and every block in $k[G]$ occurs in this way for a unique $B$ ([Isa06, Theorem 15.30(b)(g),pg. 277]). For each $B\in\mathrm{Bl}(G)$, there is a unique homomorphism $\lambda_{B}:Z(k[G])\to k$ such that $\lambda_{B}(e_{B})=1$ and $\lambda_{B}(e_{B^{\prime}})=1$ for all $B^{\prime}\neq B$ [Isa06, Theorem 15.30(a)(f),pg. 277].
Associated to a block $B\in\mathrm{Bl}(G)$ is a conjugacy class $D(B)$ of $p$-subgroups of $G$ called the defect groups of the block. Write $e_{B}\in k[G]$ as
$e_{B}=\sum_{\mathcal{C}\in\mathrm{Conj}(G)}a_{\mathcal{C}}(B)1_{\mathcal{C}}.$
A conjugacy class $\mathcal{C}$ is called a defect class of $B$ if $a_{\mathcal{C}}(B)\neq 0$ and $\lambda_{B}(1_{\mathcal{C}})\neq 0$. Every block has a defect class by [Isa06, Theorem 15.30, pg. 277]. A subgroup $H\leq G$ is called a defect group of $B$ if $H$ is the $p$-Sylow subgroup of the centralizer of an element of a defect class of $B$. The defect groups of $B$ are a conjugacy class of subgroups by [Isa06, Corollary 15.36]. If $D\in D(B)$ is a defect group of $B$, let $G:D]_{p}$ be the index of $D$ in a $p$-Sylow subgroup of $G$ (which is also the $p$-part of the integer $[G:D]$). By [Isa06, Theorem 15.41, pg. 280], this number is given by
(3.1) $G:D]_{p}=\max_{a}\{p^{a}\ :\ p^{a}\text{ divides }\dim_{K}(\rho)\text{ for all }\rho\in B\}.$
Brauer studied the relationship between blocks of $G$ and blocks of subgroups of $G$ using the Brauer homomorphism, which we now recall. Let $P\leq G$ be a $p$-subgroup and let $C_{G}(P)$ be the centralizer of $P$ in $G$. Then the $k$-linear restriction map $k[G]\to k[C_{G}(P)]$ defines a $k$-algebra homomorphism
$\mathrm{Br}_{P}:Z(k[G])\to Z(k[C_{G}(P)])$
called the Brauer homomorphism of $P$ [Isa06, Lemma 15.32, pg. 278]. If $e_{B}$ is a block of $G$, and $f$ is a block of $k[C_{G}(P)]$, then $(P,f)$ is called a $B$-Brauer pair if $f\mathrm{Br}_{p}(e_{B})=f$. Then $G$ acts on the set of $B$-Brauer pairs by conjugation, and we let $N_{G}(P,f)$ be the stabilizer of $(P,f)$ for this action.
Brauer’s Second Main Theorem ([Isa06, Theorem 15.48, pg. 284]) gives a criterion to check when $(P,f)$ is a $B$-Brauer pair in certain cases. To state it, we need the notion of generalized decomposition numbers. Suppose that $x\in G$ is an element of $p$-power order. Then for $\rho\in\mathrm{Irr}(G)$ and $\varphi\in\mathrm{IBr}(C_{G}(x))$, there are unique $d_{\rho,\varphi}^{x}\in\mathcal{O}$, called generalized decomposition numbers, such that
(3.2) $\chi_{\rho}(xy)=\sum_{\varphi\in\mathrm{IBr}(C_{G}(P))}d_{\rho,\varphi}^{x}\varphi(y)$
for all $y\in C_{G}(x)$ of prime-to-$p$ order ([Isa06, Lemma 15.47, pg. 283]).
Theorem 3.1 (Brauer’s Second Main Theorem).
Let $P\leq G$ be a $p$-subgroup and suppose that there is an element $x\in G$ of $p$-power order such that $C_{G}(x)=C_{G}(P)$. Let $B\in\mathrm{Bl}(G)$ and let $B^{\prime}\in\mathrm{Bl}(C_{G}(P))$. Suppose that $\rho\in B$ and $\varphi\in\mathrm{IBr}(C_{G}(P))$ belongs to $B^{\prime}$. If $d_{\rho,\varphi}^{x}\neq 0$, then $(P,e_{B^{\prime}})$ is a $B$-Brauer pair.
3.2. Nilpotent blocks
Recall that $G$ is called $p$-nilpotent if $G=P\ltimes H$ for a Sylow-$p$ subgroup $P$ and a subgroup $H$. It is a classical theorem of Frobenius that $G$ is $p$-nilpotent if and only if $N_{G}(P)/C_{G}(P)$ is a $p$-group for every $p$-subgroup $P\leq G$. Based on this, Broué and Puig have made the following defintion [BP80]:
Definition 3.2.
A block $B\in\mathrm{Bl}(G)$ is nilpotent if $N_{G}(P,f)/C_{G}(P)$ is a $p$-group for every $B$-Brauer pair $(P,f)$.
They show that nilpotent blocks are particularly simple (see [BP80, Theorem 1.2]). In particular, the show the following in [BP80, pg. 120].
Theorem 3.3 (Broué–Puig).
Let $B$ be a nilpotent block that has an abelian defect group $D$. Then $\tilde{e}_{B}\mathcal{O}[G]$ is Morita equivalent to $\mathcal{O}[D]$.
The statement of this theorem may seem abstract, but, in practice, the map $\tilde{e}_{B}\mathcal{O}[G]\to\mathrm{M}_{n}(\mathcal{O}[D])$ can be made completely explicit and it is not difficult to see why it is an isomorphism; we illustrate this in a particular case in Remark 4.4 below.
The theorem gives us the following application.
Lemma 3.4.
Let $M$ be a $k[G]$-module such that $\mathrm{End}_{k[G]}(M)$ is commutative and let $B$ be a nilpotent block that has cyclic defect group, and suppose that $e_{B}M\neq 0$. Let $\pi$ be a simple $e_{B}k[G]$-module. Then
$\dim_{k}\mathrm{Hom}_{k[G]}(\pi,M)=\dim_{k}\mathrm{Hom}_{k[G]}(M,\pi)=1.$
Proof.
By Theorem 3.3, $e_{B}k[G]$ is Morita-equivalent to $k[D]$, and $k[D]$ is a local monogenic algebra since $D$ is a cyclic $p$-group. Hence Lemma 2.2 gives the result. ∎
Broué and Puig [BP80, pg. 118] also give the following criteria for $B$ to be nilpotent, the first of which is clear from the definition and Frobenius’s theorem, and the second of which is based on [AB79, Proposition 4.21].
Theorem 3.5 (Alperin–Broué).
Let $B\in\mathrm{Bl}(G)$.
  1. (1)
    If $G$ is $p$-nilpotent, then $B$ is nilpotent.
  2. (2)
    If there is a $B$-Brauer pair $(D,f)$ where $D$ is a defect group of $B$ such that $N_{G}(D,f)=C_{G}(D)$ and $D$ is abelian, then $B$ is nilpotent.
In the next section, we will use the Alperin–Broué criterion (2) to show that certain blocks of Deligne–Lusztig representations are nilpotent. In Section 5, we will apply criterion (1) to show that certain representations that are used in the construction of higher-depth supercuspidal representations of $\operatorname{GL}_{n}$ are in nilpotent blocks.
4. Applications of Deligne–Lusztig theory
We use Deligne–Lusztig theory to show that certain Deligne–Lusztig characters are in nilpotent blocks with cyclic defect group. This allows us to apply the results of the previous seciton to these characters.
4.1. Deligne–Lusztig theory
We summarize some of the main results of Deligne–Lusztig theory (see [DL76] and [Ser77b]). Let $\mathcal{G}$ be a reductive group over finite field $F$ of characteristic $\ell$ with a maximal torus $\mathcal{T}\subset\mathcal{G}$. Let $G=\mathcal{G}(F)$ and $T=\mathcal{G}(F)$ and let $W=N_{G}(T)/T$ be the Weyl group of $T$.
Let $\theta:T\to K^{\times}$ be a character. For $w\in W$, let $\theta^{w}(t)=\theta(wtw^{-1})$. We say that $\theta$ is in general position if $\theta^{w}\neq\theta$ for all nontrivial $w\in W$. The following theorem summarizes the parts of Deligne–Lusztig theory that we will need.
Theorem 4.1 (Deligne–Lusztig).
For each character $\theta:T\to K^{\times}$, there is a class function
$R_{\theta}:G\to K$
that is an integer linear combination of trace characters of irreducible representations of $G$ and satisfies:
  • If $u\in G$ is unipotent, then $R_{\theta}(u)$ is an integer, independent of $\theta$, and is denoted $Q_{T,G}(u)$. The function $Q_{T,G}$ and is called the Green function of $G$ and $T$.
  • If $x\in G$ is written as $x=su=us$ for $u\in G$ unipotent and $s\in G$ semisimple, then
    (4.1) $R_{\theta}(su)=\sum_{\begin{subarray}{c}g\in G/C_{G}(s)\\ g^{-1}sg\in T\end{subarray}}Q_{gTg^{-1},C_{G}(s)}(u)\theta(g^{-1}sg).$
  • If $\theta$ is in general position, then there is a $\epsilon_{G,T}\in\{\pm 1\}$, independent of $\theta$, such that $\epsilon_{G,T}R_{\theta}$ is the trace character of an irreducible representation $\pi_{\theta}$ of $G$ of dimension $\frac{\#G}{\#T\#L}$ where $L\leq G$ is an $\ell$-Sylow subgroup. In particular, if $G$ itself is a torus and $T=G$, then $\pi_{\theta}=\theta$ and $Q_{T,G}(1)=1$.
  • If $\theta$ is in general position and $\mathcal{T}$ is not contained in a proper parabolic subgroup of $\mathcal{G}$, then $\pi_{\theta}$ is cuspidal.
Proof.
The first two points follow from [DL76, Theorem 4.2] (the integrality of $R_{\theta}(u)$ follows from [DL76, Proposition 3.3]). In the third point, the irreducibility follows from [DL76, Theorem 6.8], and the dimension formula from [DL76, Theorem 7.1]. The final point is [DL76, Theorem 8.3]. ∎
4.2. Nilpotent blocks of Deligne–Lusztig characters
Let $\ell$, $G$, and $T$ be as in the previous section. In this section, we’ll consider the case where the $p$-Sylow subgroup of $T$ is a $p$-Sylow subgroup of $G$ that contains a regular semisimple element.
Example 4.2.
Let $F=\mathbb{F}_{q}$ for $q$ a power of $\ell$ and let $n$ be the multiplicative order of $q$ modulo $p$. Let $\mathcal{G}=\operatorname{GL}_{n}$ and let $\mathcal{T}\leq\mathcal{G}$ be a torus isomorphic to $\mathrm{Res}_{\mathbb{F}_{q^{n}}/\mathbb{F}_{q}}(\mathbb{G}_{m})$. Then $T=\mathbb{F}_{q^{n}}^{\times}$ is a cyclic group whose generators are regular semisimple elements of $G$ and whose $p$-Sylow subgroup is a $p$-Sylow subgroup of $G$.
For a character $\theta:T\to\mathcal{O}^{\times}$, let ${\overline{\theta}}:T\to k^{\times}$ be the reduction modulo $\varpi\mathcal{O}$.
Theorem 4.3.
Let $\theta:T\to\mathcal{O}^{\times}$ be a character. Assume that
  • The $p$-Sylow subgroup $P$ of $T$ is a $p$-Sylow subgroup of $G$,
  • $P$ contains a regular semisimple element of $G$, and
  • $\overline{\theta}$} is in general position.
Let $\Xi$ be the set of $p$-power-order $\mathcal{O}$-valued characters of $T$. Then $\theta\xi$ is in general position for all $\xi\in\Xi$ and the collection $\{\pi_{\theta\xi}:\xi\in\Xi\}$ of irreducible representations forms a nilpotent block with defect group $P$.
Proof.
Since the characters $\theta\xi$ all reduce to $\overline{\theta}$}, they are all in general position. By Theorem 4.1, there is an irreducible representation $\pi_{\theta\xi}$ of $G$ whose trace-character $\chi_{\theta\xi}$ is $\epsilon_{G,T}R_{\theta\xi}$ for a sign $\epsilon_{G,T}\in\{\pm 1\}$ that is independent of $\xi$ and $\theta$, and whose dimension $\chi_{\theta\xi}(1)$ is $\frac{\#G}{\#T\#L}$, which is prime-to-$p$ (by the assumption that $P\leq T$ is a $p$-Sylow of $G$) and independent of $\xi$ and $\theta$.
Let
$e_{\theta\xi}=\frac{\chi_{\theta\xi}(1)}{\#G}\sum_{x\in G}\chi_{\theta\xi}(x)[x^{-1}]$
be the indempotent in $K[G]$ associated to $\pi_{\theta\xi}$, and let $\tilde{e}=\sum_{\xi\in\Xi}e_{\theta\xi}$. We claim that $\tilde{e}\in\mathcal{O}[G]$. To see this, note that, by the above-summarized properties of $\chi_{\theta\xi}$, the idempotent $\tilde{e}$ can be written as
$\tilde{e}=\epsilon_{G,T}\frac{\chi_{\theta}(1)}{\#G}\sum_{x\in G}\tilde{R}_{\theta}(x)[x^{-1}]$
where $\tilde{R}_{\theta}=\sum_{\xi\in\Xi}R_{\theta\xi}$. It’s clear then that $\tilde{e}\in\frac{1}{\#G}\mathcal{O}[G]=\frac{1}{\#P}\mathcal{O}[G]$, and to show that $\tilde{e}$ is in $\mathcal{O}[G]$, it is enough to show that $\tilde{R}_{\theta}(x)\in\#P\cdot\mathcal{O}$ for all $x\in G$. Writing $x=su=us$ with $s$ semisimple and $u$ unipotent, the formula (4.1) for $R_{\theta\xi}$ implies
$\tilde{R}_{\theta}(x)$ $=\sum_{g}Q_{gTg^{-1},C_{G}(s)}(u)\theta(g^{-1}sg)\left(\sum_{\xi\in\Xi}\xi(g^{-1}sg)\right)$
$=\begin{cases}\displaystyle\#\Xi\cdot\sum_{g}Q_{gTg^{-1},C_{G}(s)}(u)\theta(g^{-1}sg)&\text{if $s$ has prime-to-$p$ order}\\ 0&\text{otherwise}\end{cases},$
where the second equality uses the usual character-sum formula. Since $\Xi=\#P$, and the values of $\theta$ and the Green functions $Q$ are integral, this shows that $\tilde{R}_{\theta}(x)~\in~\#P\cdot\mathcal{O}$ for all $x\in G$, as required.
As we recalled in Section 3.1, the fact that $\tilde{e}$ is in $\mathcal{O}[G]$ implies that $\{\pi_{\theta\chi}:\xi\in\Xi\}$ is a union of blocks. But it is clear from the character formula (4.1) that the associated Brauer characters $\hat{\chi}_{\theta\xi}$ are independent of $\xi$, so the representations $\pi_{\theta\xi}$ are contained in the same block by definition. This shows that $B=\{\pi_{\theta\xi}:\xi\in\Xi\}$ is a block with $\tilde{e}_{B}=\tilde{e}$. Since the representations $\pi_{\theta\xi}$ in the block have prime-to-$p$ dimension, (3.1) implies that every defect group of $B$ is $p$-Sylow subgroup of $G$. Since the defect groups of a block are a conjugacy class and every $p$-Sylow subgroup of $G$ is conjugate to $P$, this implies that $P$ is a defect group of $B$.
To show that $B$ is a nilpotent block, we will find a $B$-Brauer pair and apply the Alperin–Broué criterion to it. The group $C_{G}(P)$ is $T$; let $f$ be the block of $k[T]$ that contains the character $\theta$. We claim that $(P,f)$ is an $B$-Brauer pair. To check this, we use Brauer’s Second Main Theorem (Theorem 3.1) with $x$ a regular element in $P$, $\rho=\pi_{\theta}$ and $\varphi=\hat{{\overline{\theta}}}$. Write $T=P\times S$ where $S$ is the prime-to-$p$ subgroup of $T$. The simple $k[T]$-modules are in bijection with $\hat{S}=\mathrm{Hom}(S,k^{\times})=\mathrm{Hom}(T,k^{\times})$, and, for $\eta\in\hat{S}$, its Brauer character $\hat{\eta}:S\to\mathcal{O}$ is the composition $S\xrightarrow{\eta}k^{\times}\to\mathcal{O}^{\times}$ of $\eta$ with the canonical splitting. To simplify notation, for $\eta\in\hat{S}$, we denote the generalized decomposition number $d_{\pi_{\theta},\hat{\eta}}^{x}$ by $d_{\eta}$; by Brauer’s Second Main Theorem, to show that $(P,f)$ is a $B$-Brauer pair, we need only show $d_{{\overline{\theta}}}\neq 0$. In this set up, the generalized decomposition number equation (3.2) states that
(4.2) $\epsilon_{G,T}R_{\theta}(xy)=\sum_{\eta\in\hat{S}}d_{\eta}\hat{\eta}(y)$
for all $y\in S$. Since $xy$ is regular, $C_{G}(xy)=T$ and $\{g\in G\colon g^{-1}xyg\in T\}=N_{G}(T)$, so (4.1) reads
(4.3) $R_{\theta}(xy)=\sum_{w\in W(T)}\theta^{w}(xy).$
[…]
$\sum_{\eta\in\hat{S}}{\overline{d}}_{\eta}\eta(y)=\epsilon_{G,T}\sum_{w\in W(T)}{\overline{\theta^{w}}}(y),$
where ${\overline{d}}_{\eta}$ is the image of $d_{\eta}$ in $k$. Since the characters $\overline{\theta^{w}}$} are distinct by assumption, linear independence of characters implies that ${\overline{d}}_{{\overline{\theta}}}=\epsilon_{G,T}\neq 0$, so $(P,f)$ is an $e$-Brauer pair.
Finally, $N_{G}(P,f)/C_{G}(P)$ is the set of elements in the $W(T)$ that preserve $\overline{\theta}$}, which is trivial by the assumption that $\overline{\theta}$} is in general position. Hence the hypotheses of Theorem 3.5(2) are satisfied and $e$ is a nilpotent block. ∎
Remark 4.4.
For the sake of being self-contained, we make the isomorphism $\tilde{e}_{B}\mathcal{O}[G]\cong\mathrm{M}_{n}(\mathcal{O}[P])$ of Theorem 3.3 completely explicit in this case, under the additional assumption that $P$ is cyclic, and describe the steps needed to show it is an isomorphism. First, there is the $\mathcal{O}$-algebra homomorphism $\iota:\mathcal{O}[P]\to Z(\tilde{e}_{B}\mathcal{O}[G])$ given by $\iota(g)=\sum_{\xi\in\Xi}\xi(g)e_{\theta\xi}$; a priori $\iota(g)\in Z(\tilde{e}_{B}K[G])$, but a similar argument to the proof of the theorem shows that this element is integral. If $x\in P$ is a generator, then one can calculate that
$\iota(x-1)^{|P|-1}\not\equiv 0\pmod{\varpi\mathcal{O}[G]}$,
and it follows that $\iota$ is split-injective as $\mathcal{O}$-modules. If $\mathcal{P}$ denotes a projective $\mathcal{O}[G]$-module such that $\mathcal{P}\otimes_{\mathcal{O}}k$ is a projective cover of ${\overline{\pi}}_{\theta}$, then it is easy to see that $\tilde{e}_{B}\mathcal{O}[G]\cong\mathcal{P}^{n}$ for some $n$, so that the map $Z(\tilde{e}_{B}\mathcal{O}[G])\to\mathrm{End}_{\mathcal{O}[G]}(\mathcal{P})$ is split-injective. Finally, one can show that
$\mathcal{P}\otimes_{\mathcal{O}}K\cong\bigoplus_{\xi\in\Xi}\pi_{\xi\theta}$
so that the $\mathcal{O}$-rank of $\mathrm{End}_{\mathcal{O}[G]}(\mathcal{P})$ is $\Xi=\#P$. Hence the composition of split-injections
$\mathcal{O}[P]\to Z(\tilde{e}_{B}\mathcal{O}[G])\to\mathrm{End}_{\mathcal{O}[G]}(\mathcal{P})$
is an isomorphism. This shows that
$\tilde{e}_{B}\mathcal{O}[G]\cong\mathrm{End}_{\tilde{e}_{B}\mathcal{O}[G]}(\mathcal{P}^{n})\cong\mathrm{M}_{n}(\mathrm{End}_{\mathcal{O}[G]}(\mathcal{P}))\cong\mathrm{M}_{n}(\mathcal{O}[P]).$
5. Application to $\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ and to supercuspidal representations
We illustrate how the results of the previous section may be applied in the simple example where $\mathcal{G}=\operatorname{GL}_{2}$. We also apply our results to the case of a $p$-nilpotent group, and explain how this can be used to the study of higher-depth supercuspidal representations of $\ell$-adic groups.
5.1. Cuspidal representations of $\operatorname{GL}_{2}(\mathbb{F}_{\ell})$
Let $\ell$ be a prime with $\ell\equiv-1\pmod{p}$, and let $T\cong\mathbb{F}_{\ell^{2}}^{\times}$ a non-split maximal torus in $G$. Let $Z\subset G$ be the center (the scalar matrices) and let $U\subset G$ be the upper-triangular unipotent matrices.
Let ${\overline{\theta}}:\mathbb{F}_{\ell^{2}}^{\times}\to k^{\times}$ be a character that does not factor through the norm $\mathbb{F}_{\ell^{2}}^{\times}\to\mathbb{F}_{\ell}^{\times}$, and let $\theta:\mathbb{F}_{\ell^{2}}^{\times}\to\mathcal{O}^{\times}$ be a lift of $\overline{\theta}$}. Let $\pi_{\theta}$ be an $\mathcal{O}$-lattice in the cuspidal representation of $G$ associated to $\theta$ (this lattice is unique up to homothety because the $\pi_{\theta}\otimes_{\mathcal{O}}k$ is irreducible), and let ${\overline{\pi}}_{\theta}=\pi_{\theta}\otimes_{\mathcal{O}}k$. Let $\lambda:ZU\to\mathcal{O}^{\times}$ be a character that agrees with $\theta$ on $Z$ and is non-trivial on $U$. Let $B\in\mathrm{Bl}(G)$ be the block that contains $\pi_{\theta}$ and let $\tilde{e}_{B}\in\mathcal{O}[G]$ be the associated idempotent.
Corollary 5.1.
With the above notation, there is an exact sequence of $\mathcal{O}[G]$-modules
$0\to\pi_{\theta}\to\tilde{e}_{B}\,\mathrm{Ind}_{ZU}^{G}\lambda\to\tilde{e}_{B}\,\mathrm{Ind}_{T}^{G}\theta\to 0.$
Proof.
We have to verify the hypotheses of Theorem 1.5 with $H=T$, $\lambda=\theta$, and $M_{\mathcal{O}}=\mathrm{Ind}_{ZU}^{G}\lambda$. The hypotheses on $\theta$ hold by assumption. Note that $M_{\mathcal{O}}$ is a projective $\mathcal{O}[G]$-module because $ZU$ is a prime-to-$p$ group. It remains to check that, for all $s\in G$, $\theta^{-1}\theta^{s}:T_{s}\to\mathbb{C}^{\times}$ is either trivial or does not have $p$-power order. If $s$ is in $T$, then $\theta^{-1}\theta^{s}$ is trivial, and if $s$ is not in the normalizer of $T$, then $T_{s}=Z$ is a prime-to-$p$ group, so the order of $\theta^{-1}\theta^{s}$ is prime to $p$. If $s$ is in the normalizer of $T$ but not in $T$, then the assumption that $\overline{\theta}$} is in general position implies that ${\overline{\theta}}\neq{\overline{\theta}}^{s}$, so $\theta^{-1}\theta^{s}$ does not have $p$-power order. ∎
5.2. Representations of $p$-nilpotent groups and applications to supercuspidal representations
Cuspidal representations of $\operatorname{GL}_{2}(\mathbb{F}_{\ell})$ give rise to depth-zero supercuspidal representations of $\operatorname{GL}_{2}(\mathbb{Q}_{\ell})$. To study higher-depth supercuspidal representations, instead of using Deligne–Lusztig theory, we can rely on the fact that relevant groups are $p$-nilpotent. We have the following general result for $p$-nilpotent groups.
Theorem 5.2.
Let $G$ be a finite group with a prime-to-$p$-order normal subgroup $N$ and cyclic $p$-Sylow subgroup $P$ such that $G=N\rtimes P$. Let $H\leq N$ be a subgroup and $\rho:G\to\operatorname{GL}_{n}(K)$ be a representation of $G$ such that $\rho|_{H}$ is irreducible. Let $\overline{\rho}$} be the reduction of a lattice $\rho_{\mathcal{O}}$ in $\rho$, $B\in\mathrm{Bl}(G)$ be the block that contains $\rho$, and $\tilde{e}_{B}\in\mathcal{O}[G]$ be the associated idempotent. Let $L\leq G$ be a subgroup and let $\psi:L\to\mathcal{O}^{\times}$ be a character of prime-to-$p$ order and assume that $\mathrm{Ind}_{L}^{G}\psi$ is multiplicity free.
Then there is a surjective $\mathcal{O}[G]$-module homomorphism
$\tilde{e}_{B}\mathrm{Ind}_{H}^{G}(\rho_{\mathcal{O}}|_{H})\to\tilde{e}_{B}\mathrm{Ind}_{L}^{G}\psi.$
Proof.
Since $G$ is $p$-nilpotent, $B$ is a nilpotent block, and the defect groups of $B$ are cyclic (since they are subgroups of $P$). By Proposition A.3, $\mathrm{End}_{k[G]}(\mathrm{Ind}_{L}^{G}{\overline{\psi}})$ is commutative. Thus the hypotheses of Lemma 3.4 are satisfied and
$\dim_{k}\mathrm{Hom}_{k[G]}(\mathrm{Ind}_{L}^{G}{\overline{\psi}},{\overline{\rho}})\leq 1.$
Since #H\#H is prime-to-$p$, $\tilde{e}_{B}\mathrm{Ind}_{H}^{G}(\rho_{\mathcal{O}}|_{H})$ is a projective $\mathcal{O}[G]$-module, and it is nonzero since $\mathrm{Ind}_{H}^{G}(\rho|_{H})$ contains $\rho$. The result follows from Lemma 1.1. ∎
This theorem is potentially useful when understanding the modular representation theory of positive-depth supercuspidal representations of $\operatorname{GL}_{n}$, as constructed by Howe [How77]. In the following example, we consider the case of $\operatorname{GL}_{2}$, following Bushnell–Henniart [BH06].
Example 5.3.
Let $\ell$ be a prime with $\ell\equiv-1\pmod{p}$ and let:
  • $H$ be an $\ell$-group with cyclic center $Z$, such that $V=H/Z$ is an two-dimensional $\mathbb{F}_{\ell}$-vector space,
  • $\theta:Z\hookrightarrow K^{\times}$ be a faithful character,
  • AA be a cyclic group of order $\ell+1$ that acts on $H$ in such a way that it acts on $Z$ trivially, and, for all nontrivial $a\in A$, the $a$-fixed part $V^{a}$ of $V$ is zero.
Let $G=H\rtimes A$. In this situation, there is a unique irreducible representation $\eta$ of $H$ such that $\eta|_{Z}$ contains $\theta$ [BH06, Lemma 2 of Section 16.4, pg. 114]. This representation therefore extends to a representation $\tilde{\rho}$ of $G$, but this extension is not unique. In [BH06, Lemma of Section 22.2, pg. 135], it is shown that there is a unique extension $\eta_{1}$ of $\eta$ to $G$ such that $\mathrm{tr}(\eta_{1}(a))=-1$ for all nontrivial $a\in A$, and the construction of this $\eta_{1}$ is a key step in constructing supercuspial representations of $\operatorname{GL}_{2}(\mathbb{Q}_{p})$ of positive even depth. In the course of the proof of loc. cit., the representation $\eta_{1}$ is described by a character formula
$\eta_{1}=\mathrm{Ind}_{H}^{G}\eta-\mathrm{Ind}_{ZA}^{G}\theta$
and it is shown that $\mathrm{Ind}_{ZA}^{G}\theta$ is multiplicity free. Using Theorem 5.2, we can make an analogous statement to this character formula in the context of $\mathcal{O}[G]$-modules. Indeed, writing $A=P\times A^{\prime}$, where $P$ is the $p$-Sylow subgroup of $A$, we can apply Theorem 5.2 with $N=H\rtimes A^{\prime}$, $\rho=\tilde{\eta}$, $L=ZA$, and $\psi=\theta$ to obtain an exact sequence of $\mathcal{O}[G]$-modules
(5.1) $0\to\eta_{1}\to\tilde{e}_{B}\mathrm{Ind}_{H}^{G}\eta\to\tilde{e}_{B}\mathrm{Ind}_{ZA}^{G}\theta\to 0.$
Example 5.4.
We now explain how Example 5.3 relates to even-depth supercuspidal representations of $\operatorname{GL}_{2}(F)$, where $F$ is an $\ell$-adic local field. To simplify the discussion, we only consider the case where $\ell>2$ and $F=\mathbb{Q}_{\ell}$. Let $\mathbb{Q}_{\ell^{2}}=W(\mathbb{F}_{\ell^{2}})[1/\ell]$ be the unramified quadratic extension of $\mathbb{Q}_{\ell}$, and, for $r>0$ and $i\geq 1$, let $U_{\ell^{r}}^{i}=\ker(W(\mathbb{F}_{\ell^{r}})^{\times}\to W_{i}(\mathbb{F}_{\ell^{r}})^{\times})$ and let $U_{\ell^{r}}=W(\mathbb{F}_{\ell^{r}})^{\times}$.
Let $\theta_{0}:\mathbb{Q}_{\ell^{2}}^{\times}\to K^{\times}$ be a character, let $n$ be the smallest integer such that $\theta_{0}|_{U_{\ell^{2}}^{n+1}}=1$, and assume that $n$ is positive and even. Assume that $\theta_{0}|_{U_{\ell^{2}}^{n}}$ does not factor through the norm $U_{\ell^{2}}^{n}\to U_{\ell}^{n}$. This $n$ is the depth of the corresponding supercuspidal representation. Let $\psi:\mathbb{Z}_{\ell}\to K^{\times}$ be a character with the property that $n$ is the smallest integer such that $\psi|_{\ell^{n}\mathbb{Z}_{\ell}}=1$. There is an element $u\in\mathbb{Q}_{\ell^{2}}^{\times}$ such that
$\theta_{0}(1+x)=\psi(\mathrm{tr}_{\mathbb{Q}_{\ell^{2}}/\mathbb{Q}_{\ell}}(ux))$
for all $x\in\ell^{n/2}W(\mathbb{F}_{\ell^{2}})$; fix such a $u$.
For an integer $r\geq 0$, let $\Gamma(\ell^{n})=\ker\left(\operatorname{GL}_{2}(\mathbb{Z}_{\ell})\to\operatorname{GL}_{2}(\mathbb{Z}/\ell^{r}\mathbb{Z})\right)$. Fix an embedding $U_{\ell^{2}}\subset\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$, and consider the following subgroups of $\operatorname{GL}_{2}(\mathbb{Q}_{\ell})$:
$J^{\prime\prime}=U_{\ell^{2}}^{1}\Gamma(\ell^{n/2+1})$
$J^{\prime}=U_{\ell^{2}}^{1}\Gamma(\ell^{n/2})$
$J=\mathbb{Q}_{\ell^{2}}^{\times}\Gamma(\ell^{n/2})$
Let $\theta:J^{\prime\prime}\to K^{\times}$ be the character defined by
$\theta(x(1+\ell y))=\theta_{0}(x)\psi(\mathrm{tr}(uy))$
for all $x\in U_{\ell^{2}}^{1}$ and $y\in\ell^{n/2}M_{2}(\mathbb{Z}_{\ell})$; it is well-defined by the definition of $u$. Then $\ker(\theta)$ is a normal subgroup of $J$.
We can now apply Example (5.3) to this situation. Let $H=J^{\prime}/\ker(\theta)$ and $A=\mathbb{Q}_{\ell^{2}}^{\times}/U_{\ell^{2}}^{1}\mathbb{Q}_{\ell}^{\times}\cong\mathbb{F}_{\ell^{2}}^{\times}/\mathbb{F}_{\ell}^{\times}$, which is a cyclic group of order $\ell+1$, and $G=H\rtimes A$. The center of $H$ is $Z=J^{\prime\prime}/\ker(\theta)$, on which $\theta$ is a faithful character. The hypotheses of Example (5.3) are satisfied by [BH06, Section 22.4, pg. 136]. For $R\in\{\mathcal{O},K\}$, let $\mathcal{C}_{R}(G,\theta_{0})$ denote the category of $R[G]$-modules $M$ such that $M|_{Z}$ is (multiple copies of) $\theta_{0}$. There is an exact functor
$F_{R}:\mathcal{C}_{R}(G,\theta_{0})\to R[J/\ker(\theta)]\mathrm{-mod}$
defined as follows. Let $\mathbb{Q}_{\ell^{2}}^{\times}$ act on $H$ via the quotient map $\mathbb{Q}_{\ell^{2}}^{\times}\twoheadrightarrow A$, and let $G^{\prime}=H\rtimes\mathbb{Q}_{\ell^{2}}^{\times}$, so that $G$ is a quotient of $G^{\prime}$. There is a surjective group homomorphism
$\phi:G^{\prime}\to J/\ker(\theta),\ \phi(x,h)=xh$
with $\ker(\phi)$ equal to the set of $(x,z)\in U_{\ell^{2}}^{1}\times Z$ such that $z$ is the image of $x^{-1}$. Let $\pi\in\mathcal{C}_{R}(G,\theta_{0})$, and let $\tilde{\pi}$ and $\tilde{\theta}_{0}$ be the inflations of $\pi$ and $\theta_{0}$, respectively, to $G^{\prime}$. Since the central character of $\pi$ is $\theta_{0}$, the representation $\tilde{\theta}_{0}\otimes_{R}\tilde{\pi}$ is trivial on $\ker(\phi)$, so it factors through $\phi$, giving a $R[J/\ker(\theta)]$-module. Define $F_{R}(\pi)=\tilde{\theta}_{0}\otimes_{R}\tilde{\pi}$.
Example 5.3 defines a representation $\eta_{1}$ in $\mathcal{C}_{K}(G,\theta_{0})$; the supercuspidal representation $\pi_{\theta_{0}}$ associated to $\theta_{0}$ is defined as the compact induction from $J$ to $\operatorname{GL}_{2}(\mathbb{Q}_{\ell})$ of the inflation of $F_{K}(\eta_{1})$ to $J$. The representation $F_{K}(\eta_{1})$ itself (or sometimes $\mathrm{Ind}_{J\cap\operatorname{GL}_{2}(\mathbb{Z}_{\ell})}^{\operatorname{GL}_{2}(\mathbb{Z}_{\ell})}(F_{K}(\eta_{1})|_{J\cap\operatorname{GL}_{2}(\mathbb{Z}_{\ell})})$) is called the cuspidal type of $\pi_{\theta_{0}}$. Applying $F_{\mathcal{O}}$ to (5.1) gives a modular analog of the character formula for the cuspidal type of $\pi_{\theta_{0}}$.
Appendix A Erratum to “Modular Gelfand pairs and multiplicity-free representations”
patrick b. allen, preston wake, and robin zhang
A.1. Introduction
The purpose of this appendix is to correct some errors in [Zha24]. One of the main results of that paper is [Zha24, Corollary 4.8], which claims that for a finite group $G$ with a subgroup $H$ and a complex representation $\rho$ of $H$ such that the mod-$p$ reduction $\overline{\rho}$} is irreducible, if $\mathrm{Ind}_{H}^{G}(\rho)$ is multiplicity free, then the socle and cosocle of $\mathrm{Ind}_{H}^{G}({\overline{\rho}})$ are multiplicity free. This is false in general, as the following example (which is explained fully in Section A.2.2 below) illustrates.
Example A.1.
Let $G=D_{30}$ be the dihedral group of order $30$, $H$ be the Sylow-$5$ subgroup of $H$, $\rho:H\to\mathbb{C}^{\times}$ be a non-trivial character, and let $p=5$. Then $\mathrm{Ind}_{H}^{G}(\rho)$ is multiplicity free, but both the socle and cosocle of $\mathrm{Ind}_{H}^{G}({\overline{\rho}})$ are not.
The strategy of [Zha24] is to extend the classical observation that a semisimple module $M$ is multiplicity free if and only if $\mathrm{End}(M)$ is commutative. Neither implication is true for without the semisimplicity assumption (as [Zha24, Non-examples 3.4, 3.5] illustrate). However, if the natural maps $\mathrm{End}(M)\to\mathrm{End}(\mathrm{soc}(M))$ and $\mathrm{End}(M)\to\mathrm{End}(\mathrm{cosoc}(M))$ are surjective, then this does show that the commutativity of $\mathrm{End}(M)$ implies that the semisimple modules $\mathrm{soc}(M)$ and $\mathrm{cosoc}(M)$ are multiplicity free. These two respective maps are surjective if $M$ is injective or projective, respectively, or even if $M$ satisfies the weaker condition of being self-injective or self-projective, respectively. The proof of [Zha24, Corollary 4.8] proceeds by showing that, if $\mathrm{Ind}_{H}^{G}\rho$ has commutative endomorphism ring, then so does $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ (in the proof of [Zha24, Corollary 4.8]), and that, if $\overline{\rho}$} is irreducible, then $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ is both self-injective and self-projective ([Zha24, Lemma 4.3(iii-iv)]). However, the proofs of [Zha24, Corollary 4.8] and [Zha24, Lemma 4.3(iii-iv)] contain errors, and both of these claims are false in the generality stated.
Our corrected version of [Zha24, Corollary 4.8] can be described as follows:
  • If $\mathrm{Ind}_{H}^{G}\rho$ is multiplicity free, then $\mathrm{End}_{G}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$ need not be commutative in general, but we give a necessary and sufficient criterion for this to hold (Proposition A.3). We give examples, including Example A.1, where this criterion is not satisfied.
  • If $\overline{\rho}$} is irreducible, then $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ need not be self-projective or self-injective. We prove a necessary criterion (Lemma A.12) and use it exhibit counterexmples to [Zha24, Lemma 4.3(iii-iv)]. We give simple sufficient criteria (Lemma A.14) that include:
    • $p\nmid\#H$, or
    • $H$ is normal in $G$,
    and also review a criterion due to Vignéras. Our necessary criterion is very far from these sufficient criteria and the problem of precisely determining when $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ is self-projective or self-injective seems difficult.
When the sufficient criteria for these two points are both met, then one can conclude that the socle and cosocle of $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ are multiplicity free. One simple case where the criteria are both met is when $H$ has prime-to-$p$ order and $\rho$ is a character. This case recovers [Zha24, Corollaries 5.2 and 5.4] about multiplicity-freeness of modular Gelfand–Graev modules (indeed, as noted in [Zha24, Remark 5.3], these statements are vacuous in the equal-characteristic case $\ell\mid q$, while the relevant group $H$ has prime-to-$p$ order in the unequal characteristic case $\ell\nmid q$). This also recovers the applications of [Zha24] used in [BGGG+23, Section 4.1].
A.1.1. Further comments
To avoid possible confusion about how errors in [Zha24] affect the results here, we have made this appendix completely independent of [Zha24] and do not assume that the reader is familiar with [Zha24]. We have attempted to reproduce corrected versions of the most important results of [Zha24] here, so that the reader interested in those results need only look at this paper (although we note that [Zha24, Sections 2, 3, and 6] are not affected by the errors that we identify). For the most part, we only deal with finite groups and have not attempted to extend the results to infinite compact groups, which are the subject of [Zha24, Corollary 1.10 and Sections 4.3, 5.2]. However, in Section A.4, we reproduce a proof for the $\ell=p$ case of [Zha24, Corollary 1.10] that we learned from Dipendra Prasad. We do not know whether the $\ell\neq p$ case of [Zha24, Corollary 1.10] is true because the methods we develop here are not sufficient to determine whether self-injectivity holds.
A.2. Endomorphisms of induced representations
We use Mackey theory to work out when the reduction map on twisted Hecke algebras is surjective.
A.2.1. Reduction modulo $p$ on endomorphisms of inductions
Let $\mathcal{O}$ be a discrete valuation ring with residue field $k$, uniformizer $\varpi$, and fraction field $K$. For an $\mathcal{O}$-module $M$, write ${\overline{M}}=M/\varpi M$ and $M_{K}=M\otimes_{\mathcal{O}}K$. Let $p$ be the characteristic of $k$.
Lemma A.2.
Let $R$ be an $\mathcal{O}$-algebra that is finitely generated and free as an $\mathcal{O}$-module, and let $M$ be a finitely generated $R$-module that is free as an $\mathcal{O}$-module. Then the map
${\overline{\mathrm{End}_{R}(M)}}\to\mathrm{End}_{{\overline{R}}}({\overline{M}})$
induced by $\phi\mapsto{\overline{\phi}}$ is injective. It is surjective if and only if
$\dim_{K}(\mathrm{End}_{R_{K}}(M_{K}))\geq\dim_{k}(\mathrm{End}_{{\overline{R}}}({\overline{M}})).$
Proof.
Applying the functor $\mathrm{Hom}_{R}(M,-)$ to the exact sequence
$0\to M\xrightarrow{\varpi}M\to{\overline{M}}\to 0$
yields an exact sequence
$0\to\mathrm{End}_{R}(M)\xrightarrow{\varpi}\mathrm{End}_{R}(M)\to\mathrm{Hom}_{R}(M,{\overline{M}}).$
Since $\mathrm{Hom}_{R}(M,{\overline{M}})=\mathrm{End}_{{\overline{R}}}({\overline{M}})$, this proves the first statement, and also shows that $\mathrm{End}_{R}(M)$ is a free $\mathcal{O}$-module. Since the map $\mathcal{O}\to K$ is flat, the map
$\mathrm{End}_{R}(M)_{K}\to\mathrm{End}_{R_{K}}(M_{K})$
is an isomorphism, so $\mathrm{End}_{R}(M)$ is a free $\mathcal{O}$-module of rank $\dim_{K}(\mathrm{End}_{R_{K}}(M_{K}))$. This implies that $\dim_{k}({\overline{\mathrm{End}_{R}(M)}})=\dim_{K}(\mathrm{End}_{R_{K}}(M_{K}))$, which, together with the first statement, proves the second statement. ∎
Let $G$ be a finite group, $H$ be a subgroup of $G$, and $\rho:H\to\operatorname{GL}_{n}(L)$ be a finite-dimensional representation over a field $L$. For an element $g\in H$, let $H_{g}=H\cap g^{-1}Hg$, and let $\rho^{g}$ be the representation of $H_{g}$ given by $\rho^{g}(x)=\rho(gxg^{-1})$. Let $\G/H$$\backslash G/H$$\subset$ G be a set of coset representatives. Then Mackey’s formula states that
[…]
Applying Frobenius reciprocity twice then implies that
[…]
Proposition A.3.
Let $\rho:H\to\operatorname{GL}_{n}(\mathcal{O})$ be an $\mathcal{O}$-lattice in a representation $\rho_{K}$ over $K$. Then the map
${\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}\rho)}}\to\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$
is injective and is an isomorphism if and only if
[…]
Proof.
The first statement is simply a combination of Lemma A.2 and Mackey’s formula (A.1). For $n=1$, (A.2) simplifies to
[…]
Corollary A.4.
Let $\rho:H\to\operatorname{GL}_{n}(\mathcal{O})$ be an $\mathcal{O}$-lattice in a representation $\rho_{K}$ over $K$, and suppose that $\mathrm{Ind}_{H}^{G}\rho_{K}$ is multiplicity free and that (A.2) holds. Then $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$ is commutative.
Proof.
Since $\mathrm{Ind}_{H}^{G}(\rho_{K})$ is multiplicity free, $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}\rho_{K})$ is commutative. Hence the subquotient ring $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}\rho)}$} is a also commutative. By Proposition A.3, $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}{\overline{\rho}})$ is isomorphic to $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}\rho)}$}, so it is also commutative. ∎
A.2.2. Examples
We give some examples where (A.2) fails and where the mod-$p$ twisted Hecke algebra is noncommutative even where the characteristic zero version is commutative.
Example A.5.
Let $p$ be an odd prime, $\mathcal{O}=\mathbb{Z}_{p}[\zeta_{p}]$, $G=\operatorname{GL}_{2}(\mathbb{F}_{p})$, $H$ be the subgroup of upper-triangular unipotent matrices, and $\rho:H\to\mathcal{O}^{\times}$ be a nontrivial character. By (A.1), the dimension of $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}\rho)$ is equal to
[…]
We now come to Example A.1 from the introduction. We first give a simple criterion for $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}\rho)$ to be commutative.
Lemma A.6.
Let $N$ be a normal subgroup of $G$ and let $\rho:N\to K^{\times}$ be a character. Since $N$ is normal, $\rho^{s}$ is also a character of $N$ for all $s\in G$. Let $H=\{s\in G\ |\ \rho^{s}=\rho\}$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)$ is commutative (and $\mathrm{Ind}_{N}^{G}\rho$ is multiplicity free).
Proof.
Since $N$ is normal, $\G/N$$N\backslash G$/N=G/N. By (A.1),
$\dim_{K}\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)=\#\{s\in G/N\ |\ \rho^{s}=\rho\}=[H:N].$
Since $\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)$ is a semisimple $K$-algebra, the fact that this dimension is less than $4$ implies that it is commutative. ∎
We also have a simple criterion for the socle and cosocle of $\mathrm{Ind}_{H}^{G}{\overline{\rho}}$ to have multiplicity.
Lemma A.7.
Let $G$ be a group with a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Then for every character $\rho:N\to K^{\times}$, the induction $\mathrm{Ind}_{N}^{G}{\overline{\rho}}$ is semisimple and is not multiplicity free.
Proof.
Since $N$ is a $p$-group, ${\overline{\rho}}=1$ is the trivial character. Since $N$ is normal, $\mathrm{Ind}_{N}^{G}1$ is the inflation of the regular representation $k[G/N]$. Because $p\nmid\#(G/N)$, the usual formula for the regular representation in characteristic zero,
$k[G/N]\cong\bigoplus_{\pi\in\mathrm{Irr}(G/N)}\pi^{\oplus\dim_{k}(\pi)},$
holds over $k$. Since $G/N$ is non-abelian, it has at least one irreducible representation that is not one-dimensional, so $k[G/N]$ is not multiplicity free. ∎
Combining these two lemmas gives the following.
Corollary A.8.
Let $G$ be a group that has a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Let $\rho:N\to K^{\times}$ be a character and let $H$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{Ind}_{N}^{G}\rho$ is multiplicity free but $\mathrm{Ind}_{N}^{G}{\overline{\rho}}$ is not.
A minimal example of a group satisfying the hypothesis of Corollary A.8 is $G=D_{30}$, the dihedral group of order $30$, with $p=5$ as in Example A.1. The $5$-Sylow subgroup $N$ is normal and for every nontrivial $\rho:N\to K^{\times}$, the stabilizer $H$ of $\rho$ is the centralizer $C_{G}(N)$ of $N$, which is cyclic of order $15$. Hence $[H:N]=3<4$, and Corollary A.8 applies.
A.3. Self-projectivity of induced representations
In this section, we consider the question of when an induced representation is a self-projective (or self-injective) module. Self-projectivity is a weak notion of projectivity that is enough to ensure that the map
$\mathrm{End}(M)\to\mathrm{Hom}(M,\mathrm{cosoc}(M))\cong\mathrm{End}(\mathrm{cosoc}(M))$
is surjective.
For this section, $k$ denotes a field and $R$ a finite-dimensional $k$-algebra (unital but not necessarily commutative). A typical example is $R=k[G]$ for a finite group $G$. All $R$-modules considered are assumed to be finitely-generated left $R$-modules.
A.3.1. Projective covers and injective envelopes
Recall that an $R$-module surjection $\pi:N\twoheadrightarrow M$ is called essential if there is no proper submodule $Q\subsetneq N$ such that $\pi(Q)=M$. Equivalently, $\pi$ is essential if, for every proper submodule $Q\subsetneq N$, we have $Q+\ker(\pi)\subsetneq N$. An essential surjection $\pi:P\to M$ where $P$ is a projective module is called a projective cover (we also abuse notation slightly and refer to the projective module $P$ as a projective cover of $M$).
Lemma A.9.
Let $M$ be a simple $R$-module and $P$ be a projective cover of $M$. If $N$ is a finitely generated $R$-module such that no subquotient of $N$ is isomorphic to $M$, then $\mathrm{Hom}_{R}(P,N)=0$.
Proof.
Note that the cosocle $P/\mathrm{Rad}(P)$ of $P$ is isomorphic to $M$. First consider the case that $N$ is semisimple; then the assumption is that $\mathrm{Hom}_{R}(M,N)=0$. But any homomorphism $P\to N$ factors through the cosocle $M$ of $P$, so it must be zero. Now consider general $N$ and let $f:P\to N$ be a homomorphism. By the semisimple case, the composition $P\xrightarrow{f}N\to N/\mathrm{Rad}(R)N$ is zero. Hence the image of $f$ is contained in $\mathrm{Rad}(R)N$, and the result follows by induction on the length of the radical series of $N$. ∎
Dually, a submodule $N\subseteq M$ is called essential if there is no non-zero submodule $Q\subseteq M$ such that $N\cap Q=0$. If $I$ is an injective $R$-module such that $M$ is an essential submodule of $I$, then $I$ is called an injective envelope of $M$.
A.3.2. Self-projectivity and self-injectivity
Let $M$ and $N$ be $R$-modules. Recall that an $R$-module $M$ is called $N$-projective (resp. $N$-injective) if the functor $\mathrm{Hom}_{R}(M,-)$ (resp. $\mathrm{Hom}_{R}(-,M)$) sends every short exact sequence of the form
$0\to N^{\prime}\to N\to N^{\prime\prime}\to 0$
to a short exact sequence. Of course, $M$ is projective if and only if $M$ is $N$-projective for all $N$, and similarly for injective. Similarly, $N$ is semisimple if and only if every module $M$ is $N$-projective if and only if every module $M$ is $N$-injective. A module $M$ is called self-projective (resp. self-injective) if $M$ is $M$-projective (resp. $M$-injective).
In the following lemma, the implication $(2)\Rightarrow(3)$ has been proven by Arabia [Vig98, Appendix, Proposition 7] in an arbitrary abelian category. We learned from Ofer Gabber that the converse holds for modules that have a projective cover.
Lemma A.10.
Let $M$ be a finitely generated $R$-module. The following are equivalent:
  1. (1)
    There is a projective cover $P$ of $M$ such that
    $\dim_{k}\mathrm{End}_{R}(M)=\dim_{k}\mathrm{Hom}_{R}(P,M).$
  2. (2)
    There is a projective $R$-module $P$ and a surjective homomorphism $\pi:P\twoheadrightarrow M$ such that the induced injective map
    $\mathrm{End}_{R}(M)\to\mathrm{Hom}_{R}(P,M)$
    is an isomorphism.
  3. (3)
    $M$ is self-projective.
Proof.
[…]
There is a dual lemma about self-injectivity, whose proof is almost the exact dual of the proof of Lemma A.10.
Lemma A.11.
Let $M$ be a finitely generated $R$-module. Then $M$ is self-injective if and only if there is an injective envelope $I$ of $M$ such that $\dim_{k}\mathrm{End}_{R}(M)=\dim_{k}\mathrm{Hom}_{R}(M,I).$
A.3.3. Failure of self-projectivity and self-injectivity of certain inductions
In this section, let $p$ be a prime and let $k$ be a field of characteristic $p$.
Lemma A.12.
Let $G$ be a finite group and let $H$ be a non-normal subgroup of $G$. Let $M=\mathrm{Ind}_{H}^{G}1$ be the induction of the trivial representation of $H$. Suppose that $k[G]$ is a projective cover (resp. injective envelope) of $M$. Then $M$ is not self-projective (resp. self-injective).
Proof.
Suppose that $k[G]$ is a projective cover of $M$. By Lemma A.10, to show that $M$ is not self-projective, it is enough to show that
$\dim_{k}\mathrm{End}_{k[G]}(M)<\dim_{k}\mathrm{Hom}_{k[G]}(k[G],M)$
[…]
Now suppose that $k[G]$ is an injective envelope of $M$. By Lemma A.11, to show that $M$ is not self-injective, it is enough to show that
$\dim_{k}\mathrm{End}_{k[G]}(M)<\dim_{k}\mathrm{Hom}_{k[G]}(M,k[G]).$
But $k[G]$ is self-dual, so $\mathrm{Hom}_{k[G]}(M,k[G])\cong\mathrm{Hom}_{k[G]}(k[G],M^{*})=M^{*}$ and the same arguments as in the projective case give the desired inequality. ∎
Proposition A.13.
If $G$ is a $p$-group and $H$ is a non-normal subgroup of $G$, then $\mathrm{Ind}_{H}^{G}1$ is neither self-projective nor self-injective.
Proof.
By the previous lemma, it is enough to show that $k[G]$ is a projective cover and injective envelope of $\mathrm{Ind}_{H}^{G}1$. Inducing the surjection $k[H]\twoheadrightarrow k$ shows that $k[G]$ surjects onto $\mathrm{Ind}_{H}^{G}1$, so the projective cover of $\mathrm{Ind}_{H}^{G}1$ is a direct summand of $k[G]$. Similarly, inducing the injective $k\hookrightarrow k[H]$ shows that $\mathrm{Ind}_{H}^{G}1$ can be embedded in $k[G]$, so the injective envelope of $\mathrm{Ind}_{H}^{G}1$ is a direct summand of $k[G]$. But, since $k[G]$ is local, it has no proper direct summands. ∎
A.3.4. Criteria for self-projectivity of inductions
Although the last section shows that inductions of simple modules are not self-projective in general, there are some simple criteria for when they are. We continue to assume that $p$ is a prime and $k$ is a field of characteristic $p$.
Lemma A.14.
Let $G$ be a finite group with a subgroup $H$ and let $M$ be a $k[H]$-module. Then $\mathrm{Ind}_{H}^{G}M$ is self-projective if any one of the following holds:
  1. (1)
    $M$ is $(\mathrm{Ind}_{H}^{G}M)|_{H}$-projective.
  2. (2)
    $p\nmid\#H$.
  3. (3)
    $M$ is semisimple and $H$ is normal in $G$.
Proof.
If (1) holds, then $\mathrm{Ind}_{H}^{G}M$ is self-projective by a Frobenius-reciprocity argument (see [Zha24, Corollary 4.2(ii)]). If (2) holds, then $k[H]$ is a semisimple ring, so $M$ is a projective $k[H]$-module and (1) holds. If (3) holds, then $\mathrm{Ind}_{H}^{G}M|_{H}$ is a semisimple $k[H]$-module by Mackey’s formula, so every $k[H]$-module NN is $\mathrm{Ind}_{H}^{G}M|_{H}$-projective and (1) holds again. ∎
The following criterion is a slight generalization of [Vig01, Lemma 3.1], in that we show that the assumption there that $M_{\rho}$ is a direct factor of $M|_{H}$ is unnecessary.
Lemma A.15.
Let $G$ be a finite group with a subgroup $H$. Let $\rho$ be a simple $k[H]$-module and let $M=\mathrm{Ind}_{H}^{G}\rho$. Let $M_{\rho}$ be the largest $k[H]$-submodule of $M|_{H}$ that is isomorphic to a direct sum of copies of $\rho$, and let $M^{\rho}=(M|_{H})/M_{\rho}$. If no subquotient of $M^{\rho}$ is isomorphic to $\rho$, then $M$ is self-projective.
Proof.
Let $P_{\rho}$ be a projective cover of $\rho$, and let $P=\mathrm{Ind}_{H}^{G}(P_{\rho})$. Then $P$ is a projective $k[G]$-module and there is a surjection $P\twoheadrightarrow M$. By Frobenius reciprocity,
$\mathrm{End}_{k[G]}(M)=\mathrm{Hom}_{k[H]}(\rho,M|_{H})=\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$
and
$\mathrm{Hom}_{k[G]}(P,M)=\mathrm{Hom}_{k[H]}(P_{\rho},M|_{H}).$
Since $\mathrm{Hom}_{k[H]}(P_{\rho},M^{\rho})=0$ by Lemma A.9, it follows that
$\mathrm{Hom}_{k[H]}(P_{\rho},M|_{H})=\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho}).$
But every homomorphism $P_{\rho}\to M_{\rho}$ factors through the cosocle $\rho$ of $P_{\rho}$, so $\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho})\cong\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$. Altogether, this shows that $\mathrm{End}_{k[G]}(M)$ and $\mathrm{Hom}_{k[G]}(P,M)$ are isomorphic, so $M$ is self-projective by Lemma A.10
Vignéras uses this criterion to show [Vig01, Corollary 5.2] that $\mathrm{Ind}_{P}^{G}\sigma$ is self-projective if $G=\mathbf{G}(F)$ is the $F$-rational points of a reductive group for a field $F$ with $\mathrm{char}(F)\neq\mathrm{char}(k)$, $P=MU$ is a parabolic subgroup with Levi $M$ and unipotent radical $U$, and $\sigma$ is a cuspidal representation of $M$.
A.4. Quaternion division algebras
Let $\ell$ and $p$ be two primes. Let $F$ be a $p$-adic local field with residue field $\mathbb{F}_{p^{f}}$ and let $D$ be the quaternion division algebra over $F$. The main application of [Zha24] for infinite groups is [Zha24, Corollary 5.6], which asserts the uniqueness of trilinear forms on irreducible smooth representations of $D^{\times}$ with coefficients in $k=\overline{\mathbb{F}}_{\ell}$. As a result of the errors in [Zha24] identified here, we do not know whether or not [Zha24, Corollary 5.6] is true for $\ell\neq p$, but we learned from Dipendra Prasad that the $\ell=p$ case can be recovered using results of Vignéras [Vig89], as we now explain. For the remainder of the section, we take $\ell=p$ to be an odd prime and $k={\overline{\mathbb{F}}}_{p}$.
Corollary A.16 ([Zha24, Corollary 5.6] for odd $\ell=p$).
Let $p$ be an odd prime, $k=\overline{\mathbb{F}}_{p}$, $F$ be a $p$-adic local field, and $D$ be the quaternion division algebra over $F$. If $\pi_{1},\pi_{2},\pi_{3}$ are irreducible smooth $k$-representations of $D^{\times}$, then there exists at most one (up to isomorphism) non-zero $D^{\times}$-invariant linear form on $\pi_{1}\otimes\pi_{2}\otimes\pi_{3}$ over $k$.
The proof uses the classification of irreducible smooth $k$-representations of $D^{\times}$ due to Vignéras [Vig89], which we now recall. Let $\mathcal{O}_{F}$ be the ring of integers of $F$ and $\varpi_{F}$ be a uniformizer. Let $\mathcal{O}_{D}$ be the ring of integers in $D$, let $\mathfrak{p}_{D}\subset\mathcal{O}_{D}$ be the maximal ideal, and let $U_{D}^{1}:=1+\mathfrak{p}_{D}\subset\mathcal{O}_{D}^{\times}$. Since $U_{D}^{1}$ is a normal pro-$p$ subgroup of $D^{\times}$, every irreducible smooth $k$-representation of $D^{\times}$ is inflated from the group $D^{\times}/U_{D}^{1}$ (see [Vig89, Corollaire 8.1, pg.259]). This is not a finite group, but we can arrange for such a representation to factor through a finite quotient by twisting. For an irreducible smooth $k$-representation $\pi$ of $D^{\times}$, let $\omega_{\pi}:F^{\times}\to k^{\times}$ be its central character. If $\chi:D^{\times}\to k^{\times}$ is a character such that $\chi(\varpi_{F})^{2}=\omega_{\pi}(\varpi_{F})$, then the representation $\pi\otimes\chi^{-1}$ is trivial on $\varpi_{F}$, hence it factors through the finite group $D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}$. The valuation map defines an exact sequence
$1\to\mathcal{O}_{D}^{\times}/U_{D}^{1}\to D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}\xrightarrow{\mathrm{val}}\mathbb{Z}/2\mathbb{Z}\to 1$
that splits (indeed, there is a generator $\varpi_{D}$ of $\mathfrak{p}_{D}$ such that $\varpi_{D}^{2}=\varpi_{F}$), and we note that $\mathcal{O}_{D}^{\times}/U_{D}^{1}$ is isomorphic to $\mathbb{F}_{p^{2f}}^{\times}$. Let $G=D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}$ be this dihedral group, and write it as $G\cong C\rtimes\mathopen{}\left\langle\sigma\right\rangle\mathclose{}$, where $C=\mathbb{F}_{p^{2f}}^{\times}$ and $\sigma$ acts by the Frobenius over $\mathbb{F}_{p^{f}}$. Vignéras classifies irreducible smooth $k$-representations of $D^{\times}$ using the fact that (up to twist) every such representation is inflated from $G$.
To state the classification precisely, we need more notation. Let $E/F$ be the quadratic unramified extension and fix an $F$-algebra embedding $E\hookrightarrow D$ and let $U_{E}=E^{\times}\cap U_{D}^{1}$. For a character $\psi:E^{\times}\rightarrow k^{\times}$, let $\psi^{\dagger}$ denote the character of $E^{\times}U_{D}\rightarrow k^{\times}$ such that ${\left.\kern-1.2pt\psi^{\dagger}\right|_{E^{\times}}}=\psi$ and ${\left.\kern-1.2pt\psi^{\dagger}\right|_{U_{D}}}=1$ (note that $\psi^{\dagger}$ is well defined because $U_{E}$ is a pro-$p$ group, so that $\psi|_{U_{E}}=1$). Note that since the abelianization of $D^{\times}$ is isomorphic to $F^{\times}$ [NM43], every character of $D^{\times}$ factors through the reduced norm map $\mathrm{Nm}_{D^{\times}/F^{\times}}:D^{\times}\rightarrow F^{\times}$.
Lemma A.17 ([Vig89, Proposition 9(a), p. 260], cf. [Tok15, Proposition 2.5]).
If $\pi$ is an irreducible smooth representation of $D^{\times}$, then either
  • $\pi$ is a $1$-dimensional representation isomorphic to $\chi\circ\mathrm{Nm}$ for a character $\chi:F^{\times}\rightarrow k^{\times}$.
  • $\pi$ is a $2$-dimensional representation isomorphic to $\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi^{\dagger})$ for a character $\psi:E^{\times}\rightarrow k^{\times}$ that does not factor through the norm map $\mathrm{Nm}_{E/F}:E^{\times}\rightarrow F^{\times}$. Moreover, two such representations $\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi^{\dagger})$ and $\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi^{\prime\dagger})$ are isomorphic if and only if there is $\sigma\in\mathrm{Gal}(E/F)$ such that $\psi^{\prime}=\psi^{\sigma}$.
Proof of Corollary A.16.
We will show that, for every character $\epsilon$ of $D^{\times}$, the space $\mathrm{Hom}_{D^{\times}}\mathopen{}\left(\pi_{1}\otimes\pi_{2}\otimes\pi_{3},\epsilon\right)\mathclose{}$ has dimension at most one. Since, as discussed above, each $\pi_{i}$ has a twist that is inflated from $G=D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}=C\rtimes\mathopen{}\left\langle\sigma\right\rangle\mathclose{}$, and since this proposition is invariant under twisting, we may assume that each $\pi_{i}$ is inflated from $G$. Moreover, if at least two of $\pi_{1}$, $\pi_{2}$ and $\pi_{3}$ are characters, then $\pi_{1}\otimes\pi_{2}\otimes\pi_{3}$ is irreducible and there is nothing to show, so we assume, without loss of generality, that both $\pi_{1}$ and $\pi_{2}$ have dimension greater than one. We are now working in the semisimple category of $k[G]$-modules (since $p\nmid\#G)$, and we have to show that the multiplicity of the representation $\pi_{3}^{\vee}\otimes\epsilon$ in $\pi_{1}\otimes\pi_{2}$ is at most one, where $\pi_{3}^{\vee}$ is the contragredient of $\pi_{3}$. In fact, $\pi_{1}\otimes\pi_{2}$ is multiplicity free, as we now show.
By Lemma A.17, there are characters $\psi_{1},\psi_{2}:E^{\times}\to k^{\times}$ with $\psi_{i}\neq\psi_{i}^{\sigma}$ such that $\pi_{i}=\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi_{i}^{\dagger})$. Identifying each $\pi_{i}$ with the associated representation of $G$, this is simply $\pi_{i}=\mathrm{Ind}_{C}^{G}\psi_{i}$. Then
$\pi_{1}\otimes\pi_{2}$ $\cong\mathrm{Ind}_{C}^{G}(\psi_{1})\otimes\pi_{2}$
$\cong\mathrm{Ind}_{C}^{G}\big{(}\psi_{1}\otimes\mathrm{Res}_{C}^{G}(\pi_{2})\big{)}$
$\cong\mathrm{Ind}_{C}^{G}\big{(}\psi_{1}\otimes(\psi_{2}\oplus{\psi_{2}^{\sigma}})\big{)}$
$\cong\mathrm{Ind}_{C}^{G}(\psi_{1}\psi_{2})\oplus\mathrm{Ind}_{C}^{G}(\psi_{1}{\psi_{2}^{\sigma}})$
by the push-pull formula and Mackey’s formula. Since $\psi_{i}\neq\psi_{i}^{\sigma}$, it follows that $\psi_{1}\psi_{2}\neq\psi_{1}\psi_{2}^{\sigma}$ and $\psi_{1}^{\sigma}\psi_{2}^{\sigma}\neq\psi_{1}\psi_{2}^{\sigma}$, so (using Lemma A.17 again) this is a direct sum of two distinct two-dimensional representations that do not share a common subrepresentation. In particular, $\pi_{1}\otimes\pi_{2}$ is multiplicity free.