### v2

# Corrigendum to “Modular Gelfand pairs and multiplicity-free representations”

Abstract.
Some of the multiplicity-freeness results in “Modular Gelfand pairs and multiplicity-free representations” are stated in overly broad generality. We provide counterexamples and partial corrections.
1. Introduction
The purpose of this article is to correct some errors in [Zha24]. One of the main results of that paper is [Zha24, Corollary 4.8], which claims that for a finite group $G$ with a subgroup $H$ and a complex representation $\rho$ of $H$ such that the mod-$p$ reduction $\overline{\rho}$ is irreducible, if $\mathrm{Ind}_{H}^{G}(\rho)$ is multiplicity-free, then the socle and cosocle of $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ are multiplicity-free. This is false in general, as the following example (which is explained fully in Section 2.2 below) illustrates.
Example 1.1.
Let $G=D_{30}$ be the dihedral group of order $30$, $H$ be the Sylow-$5$ subgroup of $H$, $\rho:H\to\mathbb{C}^{\times}$ be a non-trivial character, and let $p=5$. Then $\mathrm{Ind}_{H}^{G}(\rho)$ is multiplicity-free, but both the socle and cosocle of $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ are not.
The strategy of [Zha24] is to extend the classical observation that a semisimple module $M$ is multiplicity-free if and only if $\mathrm{End}(M)$ is commutative. Neither implication is true without the semisimplicity assumption (as [Zha24, Non-examples 3.4, 3.5] illustrate). However, if the natural maps $\mathrm{End}(M)\to\mathrm{End}(\mathrm{soc}(M))$ and $\mathrm{End}(M)\to\mathrm{End}(\mathrm{cosoc}(M))$ are surjective, then this does show that the commutativity of $\mathrm{End}(M)$ implies that the semisimple modules $\mathrm{soc}(M)$ and $\mathrm{cosoc}(M)$ are multiplicity-free. These two respective maps are surjective if $M$ is injective or projective, respectively, or even if $M$ satisfies the weaker condition of being self-injective or self-projective, respectively. The proof of [Zha24, Corollary 4.8] proceeds by showing that, if $\mathrm{Ind}_{H}^{G}(\rho)$ has commutative endomorphism ring, then so does $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ (in the proof of [Zha24, Corollary 4.8]), and, if $\overline{\rho}$ is irreducible, then $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ is both self-injective and self-projective ([Zha24, Lemma 4.3(iii-iv)]). However, the proofs of [Zha24, Corollary 4.8 and Lemma 4.3(iii-iv)] contain errors:
  • In the proof of [Zha24, Corollary 4.8]: the claim that if $\mathrm{End}_{G}(\mathrm{Ind}_{H}^{G}(\rho))$ has commutative endomorphism ring then $\mathrm{End}_{G}(\mathrm{Ind}_{H}^{G}(\overline{\rho}))$ is commutative is deduced from an erroneous assertion that there is always a surjection of Hecke algebras $\mathcal{H}(G,H,\rho,\mathbb{C})\twoheadrightarrow\mathcal{H}(G,H,\overline{\rho},\overline{\mathbb{F}_{p}})$.
  • In the proof of [Zha24, Lemma 4.3(iii-iv)]: the claim that if $\overline{\rho}$ is self-projective (resp. self-injective) then $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ is self-projective (resp. self-injective) is deduced from an erroneous application of Mackey’s restriction formula.
Our corrected versions can be described as follows:
  • If $\mathrm{Ind}_{H}^{G}(\rho)$ is multiplicity free, then $\mathrm{End}_{G}(\mathrm{Ind}_{H}^{G}(\overline{\rho}))$ need not be commutative in general, but we give a necessary and sufficient criterion for this to hold (Proposition 2.2). We give examples, including Example 1.1, where this criterion is not satisfied.
  • If $\overline{\rho}$ is irreducible, then $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ need not be self-projective or self-injective. We prove a necessary criterion (Lemma 3.4) and use it exhibit counterexmples to [Zha24, Lemma 4.3(iii-iv)]. We give simple sufficient criteria (Lemma 3.6) that include:
    • $p\nmid\#H$, or
    • $H$ is normal in $G$,
    and also review a criterion due to Vignéras. Our necessary criterion is very far from these sufficient criteria and the problem of precisely determining when $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ is self-projective or self-injective seems difficult.
When the sufficient criteria for these two points are both met, then one can conclude that the socle and cosocle of $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ are multiplicity free. One simple case where the criteria are both met is when $H$ has prime-to-$p$ order and $\rho$ is a character. This case recovers [Zha24, Corollaries 5.2 and 5.4] about multiplicity-freeness of modular Gelfand–Graev modules (indeed, as noted in [Zha24, Remark 5.3], these statements are vacuous in the equal-characteristic case $\ell\mid q$, while the relevant group $H$ has prime-to-$p$ order in the unequal characteristic case $\ell\nmid q$). This also recovers the applications of [Zha24] used in [BGGG+23, Section 4.1].
To avoid possible confusion about how errors in [Zha24] affect the results here, we have made this article completely independent and do not assume that the reader is familiar with [Zha24]. We have attempted to reproduce corrected versions of the most important results of [Zha24] here, so that the reader interested in those results need only look at this paper (although we note that [Zha24, Sections 2, 3, and 6] are not affected by the errors that we identify). For the most part, we only deal with finite groups and have not attempted to extend the results to infinite compact groups, which are the subject of [Zha24, Corollary 1.10 and Sections 4.3, 5.2]. However, in Section 4, we reproduce a proof for the $\ell=p$ case of [Zha24, Corollary 1.10] that we learned from Dipendra Prasad. We do not know whether the $\ell\neq p$ case of [Zha24, Corollary 1.10] is true because the methods we develop here are not sufficient to determine whether self-injectivity holds.
Acknowledgements
R. Z. is extremely grateful to his coauthors for identifying and communicating the errors and for agreeing to work together on this note. He also thanks Karol Koziol for identifying some of the errors and pointing out Example 2.4; Ofer Gabber for explaining the equivalence of self-projectivity and almost-projectivity when there are enough projectives; and Dipendra Prasad for pointing out how to recover Corollary 4.1.
P. A. acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN-2020-05915]. P. A. a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [numéro de référence RGPIN-2020-05915]. P. W. is supported by National Science Foundation CAREER Grant No. DMS-2337830. R. Z. is supported by National Science Foundation Grant No. DMS-2303280.
2. Endomorphisms of induced representations
We use Mackey theory to work out when the reduction map on twisted Hecke algebras is surjective.
2.1. Reduction modulo $p$ on endomorphisms of inductions
Let $\mathcal{O}$ be a discrete valuation ring with residue field $k$, uniformizer $\varpi$, and fraction field $K$. For an $\mathcal{O}$-module $M$, write ${\overline{M}}=M/\varpi M$ and $M_{K}=M\otimes_{\mathcal{O}}K$. Let $p$ be the characteristic of $k$.
Lemma 2.1.
Let $R$ be an $\mathcal{O}$-algebra that is finitely generated and free as an $\mathcal{O}$-module, and let $M$ be a finitely generated $R$-module that is free as an $\mathcal{O}$-module. Then the map
${\overline{\mathrm{End}_{R}(M)}}\to\mathrm{End}_{{\overline{R}}}({\overline{M}})$
induced by $\phi\mapsto{\overline{\phi}}$ is injective. It is surjective if and only if
$\dim_{K}(\mathrm{End}_{R_{K}}(M_{K}))\geq\dim_{k}(\mathrm{End}_{{\overline{R}}}({\overline{M}})).$
Proof.
Applying the functor $\mathrm{Hom}_{R}(M,-)$ to the exact sequence
$0\to M\xrightarrow{\varpi}M\to{\overline{M}}\to 0$
yields an exact sequence
$0\to\mathrm{End}_{R}(M)\xrightarrow{\varpi}\mathrm{End}_{R}(M)\to\mathrm{Hom}_{R}(M,{\overline{M}}).$
Since $\mathrm{Hom}_{R}(M,{\overline{M}})=\mathrm{End}_{{\overline{R}}}({\overline{M}})$, this proves the first statement, and also shows that $\mathrm{End}_{R}(M)$ is a free $\mathcal{O}$-module. Since the map $\mathcal{O}\to K$ is flat, the map
$\mathrm{End}_{R}(M)_{K}\to\mathrm{End}_{R_{K}}(M_{K})$
is an isomorphism, so $\mathrm{End}_{R}(M)$ is a free $\mathcal{O}$-module of rank $\dim_{K}(\mathrm{End}_{R_{K}}(M_{K}))$. This implies that $\dim_{k}({\overline{\mathrm{End}_{R}(M)}})=\dim_{K}(\mathrm{End}_{R_{K}}(M_{K}))$, which, together with the first statement, proves the second statement. ∎
Let $G$ be a finite group, $H$ be a subgroup of $G$, and $\rho:H\to\operatorname{GL}_{n}(L)$ be a finite-dimensional representation over a field $L$. For an element $g\in H$, let $H_{g}=H\cap g^{-1}Hg$, and let $\rho^{g}$ be the representation of $H_{g}$ given by $\rho^{g}(x)=\rho(gxg^{-1})$. Let $\G/H$$\backslash G/H$$\subset$ G be a set of coset representatives. Then Mackey’s formula states that
[…]
Applying Frobenius reciprocity twice then implies that
[…]
Proposition 2.2.
Let $\rho:H\to\operatorname{GL}_{n}(\mathcal{O})$ be an $\mathcal{O}$-lattice in a representation $\rho_{K}$ over $K$. Then the map
${\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}(\rho))}}\to\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}({\overline{\rho}}))$
is injective and is an isomorphism if and only if
[…]
Proof.
The first statement is simply a combination of Lemma 2.1 and Mackey’s formula (2.1). For $n=1$, (2.2) simplifies to
[…]
Corollary 2.3.
Let $\rho:H\to\operatorname{GL}_{n}(\mathcal{O})$ be an $\mathcal{O}$-lattice in a representation $\rho_{K}$ over $K$, and suppose that $\mathrm{Ind}_{H}^{G}(\rho_{K})$ is multiplicity-free and that (2.2) holds. Then $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}(\overline{\rho}))$ is commutative.
Proof.
Since $\mathrm{Ind}_{H}^{G}(\rho_{K})$ is multiplicity-free, $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}(\rho_{K}))$ is commutative. Hence the subquotient ring $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}(\rho))}$} is a also commutative. By Proposition 2.2, $\mathrm{End}_{k[G]}(\mathrm{Ind}_{H}^{G}(\overline{\rho}))$ is isomorphic to $\overline{\mathrm{End}_{\mathcal{O}[G]}(\mathrm{Ind}_{H}^{G}(\rho))}$}, so it is also commutative. ∎
2.2. Examples
We give some examples where (2.2) fails and where the mod-$p$ twisted Hecke algebra is noncommutative even where the characteristic zero version is commutative.
Example 2.4.
Let $p$ be an odd prime, $\mathcal{O}=\mathbb{Z}_{p}[\zeta_{p}]$, $G=\operatorname{GL}_{2}(\mathbb{F}_{p})$, $H$ be the subgroup of upper-triangular unipotent matrices, and $\rho:H\to\mathcal{O}^{\times}$ be a nontrivial character. By (2.1), the dimension of $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}(\rho))$ is equal to
[…]
We now come to Example 1.1 from the introduction. We first give a simple criterion for $\mathrm{End}_{K[G]}(\mathrm{Ind}_{H}^{G}(\rho))$ to be commutative.
Lemma 2.5.
Let $N$ be a normal subgroup of $G$ and let $\rho:N\to K^{\times}$ be a character. Since $N$ is normal, $\rho^{s}$ is also a character of $N$ for all $s\in G$. Let $H=\{s\in G\ |\ \rho^{s}=\rho\}$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)$ is commutative (and $\mathrm{Ind}_{N}^{G}(\rho)$ is multiplicity-free).
Proof.
Since $N$ is normal, $\G/N$$N\backslash G$/N=G/N. By (2.1),
$\dim_{K}\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)=\#\{s\in G/N\ |\ \rho^{s}=\rho\}=[H:N].$
Since $\mathrm{End}_{K[G]}\mathrm{Ind}_{N}^{G}(\rho)$ is a semisimple $K$-algebra, the fact that this dimension is less than $4$ implies that it is commutative. ∎
We also have a simple criterion for the socle and cosocle of $\mathrm{Ind}_{H}^{G}(\overline{\rho})$ to have multiplicity.
Lemma 2.6.
Let $G$ be a group with a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Then for every character $\rho:N\to K^{\times}$, the induction $\mathrm{Ind}_{N}^{G}(\overline{\rho})$ is semisimple and is not multiplicity free.
Proof.
Since $N$ is a $p$-group, $\overline{\rho}=1$ is the trivial character. Since $N$ is normal, $\mathrm{Ind}_{N}^{G}(\mathbbm{1})$ is the inflation of the regular representation $k[G/N]$. Because $p\nmid\#(G/N)$, the usual formula for the regular representation in characteristic zero,
$k[G/N]\cong\bigoplus_{\pi\in\mathrm{Irr}(G/N)}\pi^{\oplus\dim_{k}(\pi)},$
holds over $k$. Since $G/N$ is non-abelian, it has at least one irreducible representation that is not one-dimensional, so $k[G/N]$ is not multiplicity free. ∎
Combining these two lemmas gives the following.
Corollary 2.7.
Let $G$ be a group that has a normal $p$-Sylow subgroup $N$ such that $G/N$ is non-abelian. Let $\rho:N\to K^{\times}$ be a character and let $H$ be the stabilizer of $\rho$. If $[H:N]<4$, then $\mathrm{Ind}_{N}^{G}(\rho)$ is multiplicity-free but $\mathrm{Ind}_{N}^{G}(\overline{\rho})$ is not.
A minimal example of a group satisfying the hypothesis of Corollary 2.7 is $G=D_{30}$, the dihedral group of order $30$, with $p=5$ as in Example 1.1. The 55-Sylow subgroup NN is normal and for every nontrivial $\rho:N\to K^{\times}$, the stabilizer $H$ of $\rho$ is the centralizer $C_{G}(N)$ of $N$, which is cyclic of order $15$. Hence $[H:N]=3<4$, and Corollary 2.7 applies.
3. Self-projectivity of induced representations
In this section, we consider the question of when an induced representation is a self-projective (or self-injective) module. Self-projectivity is a weak notion of projectivity that is enough to ensure that the map
$\mathrm{End}(M)\to\mathrm{Hom}(M,\mathrm{cosoc}(M))\cong\mathrm{End}(\mathrm{cosoc}(M))$
is surjective.
For this section, $k$ denotes a field and $R$ a finite-dimensional $k$-algebra (unital but not necessarily commutative). A typical example is $R=k[G]$ for a finite group $G$. All $R$-modules considered are assumed to be finitely-generated left $R$-modules.
3.1. Projective covers and injective envelopes
Recall that an $R$-module surjection $\pi:N\twoheadrightarrow M$ is called essential if there is no proper submodule $Q\subsetneq N$ such that $\pi(Q)=M$. Equivalently, $\pi$ is essential if, for every proper submodule $Q\subsetneq N$, we have $Q+\ker(\pi)\subsetneq N$. An essential surjection $\pi:P\to M$ where $P$ is a projective module is called a projective cover (we also abuse notation slightly and refer to the projective module $P$ as a projective cover of $M$).
Lemma 3.1.
Let $M$ be a simple $R$-module and $P$ be a projective cover of $M$. If $N$ is a finitely generated $R$-module such that no subquotient of $N$ is isomorphic to $M$, then $\mathrm{Hom}_{R}(P,N)=0$.
Proof.
Note that the cosocle $P/\mathrm{Rad}(P)$ of $P$ is isomorphic to $M$. First consider the case that $N$ is semisimple; then the assumption is that $\mathrm{Hom}_{R}(M,N)=0$. But any homomorphism $P\to N$ factors through the cosocle $M$ of $P$, so it must be zero. Now consider general $N$ and let $f:P\to N$ be a homomorphism. By the semisimple case, the composition $P\xrightarrow{f}N\to N/\mathrm{Rad}(R)N$ is zero. Hence the image of $f$ is contained in $\mathrm{Rad}(R)N$, and the result follows by induction on the length of the radical series of $N$. ∎
Dually, a submodule $N\subseteq M$ is called essential if there is no non-zero submodule $Q\subseteq M$ such that $N\cap Q=0$. If $I$ is an injective $R$-module such that $M$ is an essential submodule of $I$, then $I$ is called an injective envelope of $M$.
3.2. Self-projectivity and self-injectivity
Let $M$ and $N$ be $R$-modules. Recall that an $R$-module $M$ is called $N$-projective (resp. $N$-injective) if the functor $\mathrm{Hom}_{R}(M,-)$ (resp. $\mathrm{Hom}_{R}(-,M)$) sends every short exact sequence of the form
$0\to N^{\prime}\to N\to N^{\prime\prime}\to 0$
to a short exact sequence. Of course, $M$ is projective if and only if $M$ is $N$-projective for all $N$, and similarly for injective. Similarly, $N$ is semisimple if and only if every module $M$ is $N$-projective if and only if every module $M$ is $N$-injective. A module $M$ is called self-projective (resp. self-injective) if $M$ is $M$-projective (resp. $M$-injective).
In the following lemma, the implication $(2)\Rightarrow(3)$ has been proven by Arabia [Vig98, Appendix, Proposition 7] in an arbitrary abelian category. We learned from Ofer Gabber that the converse holds for modules that have a projective cover.
Lemma 3.2.
Let $M$ be a finitely generated $R$-module. The following are equivalent:
  1. (1)
    There is a projective cover $P$ of $M$ such that
    $\dim_{k}\mathrm{End}_{R}(M)=\dim_{k}\mathrm{Hom}_{R}(P,M).$
  2. (2)
    There is a projective $R$-module $P$ and a surjective homomorphism $\pi:P\twoheadrightarrow M$ such that the induced injective map
    $\mathrm{End}_{R}(M)\to\mathrm{Hom}_{R}(P,M)$
    is an isomorphism.
  3. (3)
    $M$ is self-projective.
Proof.
[…]
There is a dual lemma about self-injectivity, whose proof is almost the exact dual of the proof of Lemma 3.2.
Lemma 3.3.
Let $M$ be a finitely generated $R$-module. Then $M$ is self-injective if and only if there is an injective envelope $I$ of $M$ such that $\dim_{k}\mathrm{End}_{R}(M)=\dim_{k}\mathrm{Hom}_{R}(M,I).$
3.3. Failure of self-projectivity and self-injectivity of certain inductions
In this section, let $p$ be a prime and let $k$ be a field of characteristic $p$.
Lemma 3.4.
Let $G$ be a finite group and let $H$ be a non-normal subgroup of $G$. Let $M=\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ be the induction of the trivial representation of $H$. Suppose that $k[G]$ is a projective cover (resp. injective envelope) of $M$. Then $M$ is not self-projective (resp. self-injective).
Proof.
Suppose that $k[G]$ is a projective cover of $M$. By Lemma 3.2, to show that $M$ is not self-projective, it is enough to show that
$\dim_{k}\mathrm{End}_{k[G]}(M)<\dim_{k}\mathrm{Hom}_{k[G]}(k[G],M)$
[…]
Now suppose that $k[G]$ is an injective envelope of $M$. By Lemma 3.3, to show that $M$ is not self-injective, it is enough to show that
$\dim_{k}\mathrm{End}_{k[G]}(M)<\dim_{k}\mathrm{Hom}_{k[G]}(M,k[G]).$
But $k[G]$ is self-dual, so $\mathrm{Hom}_{k[G]}(M,k[G])\cong\mathrm{Hom}_{k[G]}(k[G],M^{*})=M^{*}$ and the same arguments as in the projective case give the desired inequality. ∎
Proposition 3.5.
If $G$ is a $p$-group and $H$ is a non-normal subgroup of $G$, then $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ is neither self-projective nor self-injective.
Proof.
By the previous lemma, it is enough to show that $k[G]$ is a projective cover and injective envelope of $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$. Inducing the surjection $k[H]\twoheadrightarrow k$ shows that $k[G]$ surjects onto $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$, so the projective cover of $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ is a direct summand of $k[G]$. Similarly, inducing the injective $k\hookrightarrow k[H]$ shows that $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ can be embedded in $k[G]$, so the injective envelope of $\mathrm{Ind}_{H}^{G}(\mathbbm{1})$ is a direct summand of $k[G]$. But, since $k[G]$ is local, it has no proper direct summands. ∎
3.4. Criteria for self-projectivity of inductions
Although the last section shows that inductions of simple modules are not self-projective in general, there are some simple criteria for when they are. We continue to assume that $p$ is a prime and $k$ is a field of characteristic $p$.
Lemma 3.6.
Let $G$ be a finite group with a subgroup $H$ and let $M$ be a $k[H]$-module. Then $\mathrm{Ind}_{H}^{G}(M)$ is self-projective if any one of the following holds:
  1. (1)
    $M$ is $\mathrm{Res}_{H}^{G}(\mathrm{Ind}_{H}^{G}(M))$-projective.
  2. (2)
    $p\nmid\#H$.
  3. (3)
    $M$ is semisimple and $H$ is normal in $G$.
Proof.
If (1) holds, then $\mathrm{Ind}_{H}^{G}(M)$ is self-projective by a Frobenius-reciprocity argument (see [Zha24, Corollary 4.2(ii)]). If (2) holds, then $k[H]$ is a semisimple ring, so $M$ is a projective $k[H]$-module and (1) holds. If (3) holds, then $\mathrm{Res}_{H}^{G}(\mathrm{Ind}_{H}^{G}(M))$ is a semisimple $k[H]$-module by Mackey’s formula, so every $k[H]$-module NN is $\mathrm{Res}_{H}^{G}(\mathrm{Ind}_{H}^{G}(M))$-projective and (1) holds again. ∎
The following criterion is a slight generalization of [Vig01, Lemma 3.1], in that we show that the assumption there that $M_{\rho}$ is a direct factor of $\mathrm{Res}_{H}^{G}(M)$ is unnecessary.
Lemma 3.7.
Let $G$ be a finite group with a subgroup $H$. Let $\rho$ be a simple $k[H]$-module and let $M=\mathrm{Ind}_{H}^{G}(\rho)$. Let $M_{\rho}$ be the largest $k[H]$-submodule of $\mathrm{Res}_{H}^{G}(M)$ that is isomorphic to a direct sum of copies of $\rho$, and let $M^{\rho}=\mathrm{Res}_{H}^{G}(M)/M_{\rho}$. If no subquotient of $M^{\rho}$ is isomorphic to $\rho$, then $M$ is self-projective.
Proof.
Let $P_{\rho}$ be a projective cover of $\rho$, and let $P=\mathrm{Ind}_{H}^{G}(P_{\rho})$. Then $P$ is a projective $k[G]$-module and there is a surjection $P\twoheadrightarrow M$. By Frobenius reciprocity,
$\mathrm{End}_{k[G]}(M)=\mathrm{Hom}_{k[H]}(\rho,\mathrm{Res}_{H}^{G}(M))=\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$
and
$\mathrm{Hom}_{k[G]}(P,M)=\mathrm{Hom}_{k[H]}(P_{\rho},\mathrm{Res}_{H}^{G}(M)).$
Since $\mathrm{Hom}_{k[H]}(P_{\rho},M^{\rho})=0$ by Lemma 3.1, it follows that
$\mathrm{Hom}_{k[H]}(P_{\rho},\mathrm{Res}_{H}^{G}(M))=\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho}).$
But every homomorphism $P_{\rho}\to M_{\rho}$ factors through the cosocle $\rho$ of $P_{\rho}$, so $\mathrm{Hom}_{k[H]}(P_{\rho},M_{\rho})\cong\mathrm{Hom}_{k[H]}(\rho,M_{\rho})$. Altogether, this shows that $\mathrm{End}_{k[G]}(M)$ and $\mathrm{Hom}_{k[G]}(P,M)$ are isomorphic, so $M$ is self-projective by Lemma 3.2
Vignéras uses this criterion to show [Vig01, Corollary 5.2] that $\mathrm{Ind}_{P}^{G}(\sigma)$ is self-projective if $G=\mathbf{G}(F)$ is the $F$-rational points of a reductive group for a field $F$ with $\mathrm{char}(F)\neq\mathrm{char}(k)$, $P=MU$ is a parabolic subgroup with Levi $M$ and unipotent radical $U$, and $\sigma$ is a cuspidal representation of $M$.
4. Quaternion division algebras
Let $\ell$ and $p$ be two primes. Let $F$ be a $p$-adic local field with residue field $\mathbb{F}_{p^{f}}$ and let $D$ be the quaternion division algebra over $F$. The main application of [Zha24] for infinite groups is [Zha24, Corollary 5.6], which asserts the uniqueness of trilinear forms on irreducible smooth representations of $D^{\times}$ with coefficients in $k=\overline{\mathbb{F}}_{\ell}$. As a result of the errors in [Zha24] identified here, we do not know whether or not [Zha24, Corollary 5.6] is true for $\ell\neq p$, but we learned from Dipendra Prasad that the $\ell=p$ case can be recovered using results of Vignéras [Vig89], as we now explain. For the remainder of the section, we take $\ell=p$ to be an odd prime and $k={\overline{\mathbb{F}}}_{p}$.
Corollary 4.1 ([Zha24, Corollary 5.6] for odd $\ell=p$).
Let $p$ be an odd prime, $k=\overline{\mathbb{F}}_{p}$, $F$ be a $p$-adic local field, and $D$ be the quaternion division algebra over $F$. If $\pi_{1},\pi_{2},\pi_{3}$ are irreducible smooth $k$-representations of $D^{\times}$, then there exists at most one (up to isomorphism) non-zero $D^{\times}$-invariant linear form on $\pi_{1}\otimes\pi_{2}\otimes\pi_{3}$ over $k$.
The proof uses the classification of irreducible smooth $k$-representations of $D^{\times}$ due to Vignéras [Vig89], which we now recall. Let $\mathcal{O}_{F}$ be the ring of integers of $F$ and $\varpi_{F}$ be a uniformizer. Let $\mathcal{O}_{D}$ be the ring of integers in $D$, let $\mathfrak{p}_{D}\subset\mathcal{O}_{D}$ be the maximal ideal, and let $U_{D}^{1}:=1+\mathfrak{p}_{D}\subset\mathcal{O}_{D}^{\times}$. Since $U_{D}^{1}$ is a normal pro-$p$ subgroup of $D^{\times}$, every irreducible smooth $k$-representation of $D^{\times}$ is inflated from the group $D^{\times}/U_{D}^{1}$ (see [Vig89, Corollaire 8.1, pg.259]). This is not a finite group, but we can arrange for such a representation to factor through a finite quotient by twisting. For an irreducible smooth $k$-representation $\pi$ of $D^{\times}$, let $\omega_{\pi}:F^{\times}\to k^{\times}$ be its central character. If $\chi:D^{\times}\to k^{\times}$ is a character such that $\chi(\varpi_{F})^{2}=\omega_{\pi}(\varpi_{F})$, then the representation $\pi\otimes\chi^{-1}$ is trivial on $\varpi_{F}$, hence it factors through the finite group $D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}$. The valuation map defines an exact sequence
$1\to\mathcal{O}_{D}^{\times}/U_{D}^{1}\to D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}\xrightarrow{\mathrm{val}}\mathbb{Z}/2\mathbb{Z}\to 1$
that splits (indeed, there is a generator $\varpi_{D}$ of $\mathfrak{p}_{D}$ such that $\varpi_{D}^{2}=\varpi_{F}$), and we note that $\mathcal{O}_{D}^{\times}/U_{D}^{1}$ is isomorphic to $\mathbb{F}_{p^{2f}}^{\times}$. Let $G=D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}$ be this dihedral group, and write it as $G\cong C\rtimes\mathopen{}\left\langle\sigma\right\rangle\mathclose{}$, where $C=\mathbb{F}_{p^{2f}}^{\times}$ and $\sigma$ acts by the Frobenius over $\mathbb{F}_{p^{f}}$. Vignéras classifies irreducible smooth $k$-representations of $D^{\times}$ using the fact that (up to twist) every such representation is inflated from $G$.
To state the classification precisely, we need more notation. Let $E/F$ be the quadratic unramified extension and fix an $F$-algebra embedding $E\hookrightarrow D$ and let $U_{E}=E^{\times}\cap U_{D}^{1}$. For a character $\psi:E^{\times}\rightarrow k^{\times}$, let $\psi^{\dagger}$ denote the character of $E^{\times}U_{D}\rightarrow k^{\times}$ such that ${\left.\kern-1.2pt\psi^{\dagger}\right|_{E^{\times}}}=\psi$ and ${\left.\kern-1.2pt\psi^{\dagger}\right|_{U_{D}}}=1$ (note that $\psi^{\dagger}$ is well defined because $U_{E}$ is a pro-$p$ group, so that $\psi|_{U_{E}}=1$). Note that since the abelianization of $D^{\times}$ is isomorphic to $F^{\times}$ [NM43], every character of $D^{\times}$ factors through the reduced norm map $\mathrm{Nm}_{D^{\times}/F^{\times}}:D^{\times}\rightarrow F^{\times}$.
Lemma 4.2 ([Vig89, Proposition 9(a), p. 260], cf. [Tok15, Proposition 2.5]).
If $\pi$ is an irreducible smooth representation of $D^{\times}$, then either
  • $\pi$ is a $1$-dimensional representation isomorphic to $\chi\circ\mathrm{Nm}$ for a character $\chi:F^{\times}\rightarrow k^{\times}$.
  • $\pi$ is a $2$-dimensional representation isomorphic to $\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi^{\dagger})$ for a character $\psi:E^{\times}\rightarrow k^{\times}$ that does not factor through the norm map $\mathrm{Nm}_{E/F}:E^{\times}\rightarrow F^{\times}$. Moreover, two such representations $\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi^{\dagger})$ and $\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi^{\prime\dagger})$ are isomorphic if and only if there is $\sigma\in\mathrm{Gal}(E/F)$ such that $\psi^{\prime}=\psi^{\sigma}$.
Proof of Corollary 4.1.
We will prove the following stronger statement: for every character $\epsilon$ of $D^{\times}$, the space $\mathrm{Hom}_{D^{\times}}\mathopen{}\left(\pi_{1}\otimes\pi_{2}\otimes\pi_{3},\epsilon\right)\mathclose{}$ has dimension at most one. Since, as discussed above, each $\pi_{i}$ has a twist that is inflated from $G=D^{\times}/\varpi_{F}^{\mathbb{Z}}U_{D}^{1}=C\rtimes\mathopen{}\left\langle\sigma\right\rangle\mathclose{}$, and since this statement is invariant under twisting, we may assume that each $\pi_{i}$ is inflated from $G$. Moreover, if at least two of $\pi_{1}$, $\pi_{2}$ and $\pi_{3}$ are characters, then $\pi_{1}\otimes\pi_{2}\otimes\pi_{3}$ is irreducible and there is nothing to show, so we assume, without loss of generality, that both $\pi_{1}$ and $\pi_{2}$ have dimension greater than one. We are now working in the semisimple category of $k[G]$-modules (since $p\nmid\#G)$, and we have to show that the multiplicity of the representation $\pi_{3}^{\vee}\otimes\epsilon$ in $\pi_{1}\otimes\pi_{2}$ is at most one, where $\pi_{3}^{\vee}$ is the contragredient of $\pi_{3}$. In fact, $\pi_{1}\otimes\pi_{2}$ is multiplicity-free, as we now show.
By Lemma 4.2, there are characters $\psi_{1},\psi_{2}:E^{\times}\to k^{\times}$ with $\psi_{i}\neq\psi_{i}^{\sigma}$ such that $\pi_{i}=\mathrm{Ind}_{E^{\times}U_{D}^{1}}^{D^{\times}}(\psi_{i}^{\dagger})$. Identifying each $\pi_{i}$ with the associated representation of $G$, this is simply $\pi_{i}=\mathrm{Ind}_{C}^{G}(\psi_{i})$. Then
$\pi_{1}\otimes\pi_{2}$ $\cong\mathrm{Ind}_{C}^{G}(\psi_{1})\otimes\pi_{2}$
$\cong\mathrm{Ind}_{C}^{G}\big(\psi_{1}\otimes\mathrm{Res}_{C}^{G}(\pi_{2})\big)$
$\cong\mathrm{Ind}_{C}^{G}\big(\psi_{1}\otimes(\psi_{2}\oplus{\psi_{2}^{\sigma}})\big)$
$\cong\mathrm{Ind}_{C}^{G}(\psi_{1}\psi_{2})\oplus\mathrm{Ind}_{C}^{G}(\psi_{1}{\psi_{2}^{\sigma}})$
by the push-pull formula and Mackey’s formula. Since $\psi_{i}\neq\psi_{i}^{\sigma}$, it follows that $\psi_{1}\psi_{2}\neq\psi_{1}\psi_{2}^{\sigma}$ and $\psi_{1}^{\sigma}\psi_{2}^{\sigma}\neq\psi_{1}\psi_{2}^{\sigma}$, so (using Lemma 4.2 again) this is a direct sum of two distinct two-dimensional representations that do not share a common subrepresentation. In particular, $\pi_{1}\otimes\pi_{2}$ is multiplicity-free. ∎